### v1

# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

Abstract.
Generative retrieval methods utilize generative sequential modeling techniques, such as transformers, to generate candidate items for recommender systems. These methods have demonstrated promising results in academic benchmarks, surpassing traditional retrieval models like two-tower architectures. However, current generative retrieval methods lack the scalability required for industrial recommender systems, and they are insufficiently flexible to satisfy the multiple metric requirements of modern systems.
This paper introduces PinRec, a novel generative retrieval model developed for applications at Pinterest. PinRec utilizes outcome-conditioned generation, enabling modelers to specify how to balance various outcome metrics, such as the number of saves and clicks, to effectively align with business goals and user exploration. Additionally, PinRec incorporates multi-token generation to enhance output diversity while optimizing generation. Our experiments demonstrate that PinRec can successfully balance performance, diversity, and efficiency, delivering a significant positive impact to users using generative models. This paper marks a significant milestone in generative retrieval, as it presents, to our knowledge, the first rigorous study on implementing generative retrieval at the scale of Pinterest.
generative retrieval, recommender systems, transformer, item recommendation
journalyear: 2025copyright: ccccs: Information systems Retrieval models and rankingccs: Information systems Personalization
Figure 1. Illustration of PinRec, a generative item retrieval technique for heterogeneous user journeys on Pinterest. Sequences of user searches, engagements, and outcome-conditioning (bottom) are used to recommend Pins (top).
1. Introduction
Pinterest is a visual discovery platform where over 550 million monthly active users search for, save, and shop for ideas, represented by Pins (Pinterest, 2024a). Given the massive volume of Pins — for instance, users save over 1.5 billion Pins per week (Pinterest, 2024b) — it is essential to provide personalized recommendations in Pinterest feeds and search results. To achieve this, recommender systems often leverage a two-stage approach: retrieval and ranking. Retrieval models select thousands of potentially relevant candidate items, then ranking models prioritize the top items to display to the user. Since ranking models depend on the retrieval results, it is important for retrieval models to identify diverse, fresh, and relevant items.
Traditionally, retrieval methods employ “two-tower” models. The query tower computes query embeddings, while the item tower computes item embeddings. These models retrieve items with the highest similarity scores given a query embedding. As an alternative to two-tower models, recent work has proposed generative retrieval. Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (Vaswani et al., 2017), and then use those tokens to retrieve items. Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018; Li et al., 2023; Yang et al., 2024). Nevertheless, to our knowledge, generative retrieval has not been attempted as successfully for large-scale industrial applications yet.
Using generative retrieval models for large scale recommender systems poses many challenges at the scale of Pinterest. Given their increased complexity and sequential generation, generating a set of recommended items requires higher computational cost than computing an embedding with two-tower models. Further, the generative model should be capable of efficiently optimizing for different engagement outcomes because (a) business needs are constantly evolving to optimize combinations of different metrics and (b) as importantly, we wish for users to engage with content in the most optimal way for their growth on the platform. For instance, we would want to approach recommendations for a user with a propensity for low-effort engagement differently than for core users; through which recommendations could we enable the maximum progress in their user journey on Pinterest? To answer this question requires a malleable technique that can adapt to different circumstances, goals, and demands. Finally, the generated item sequences must possess sufficient diversity to facilitate users’ exploration of varied ideas; one-step generation of a single item representation may not suffice.
To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges. Most methods developed within academic settings (Kang and McAuley, 2018; Hou et al., 2022; Li et al., 2023) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic. While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (Zhai et al., 2024) and do not demonstrate optimal optimization for different target outcomes.
In this paper, we present PinRec, a generative retrieval technique for real-world recommender systems. Unlike previous methods, PinRec addresses and resolves three crucial criteria for generation of optimal recommendations: multifaceted generative capabilities, computational efficiency, and item diversity. Our main contributions in the paper are summarized below.
  • Outcome-Conditioned Generative Retrieval: Given evolving business needs and to allow for different types of user exploration, we propose multifaceted generative capabilities through outcome-conditioned generation. Across offline and online evaluation, we demonstrate that customizing item recommendations to the desired outcome allows for designers to provide recommendations that successfully result in those specified outcomes.
  • Windowed, Multi-Token Prediction: We propose an efficient multi-token prediction objective tailored to the structure of the sequential recommendation task, where prior next-token objectives have often lacked representational power. Instead, by optimizing multiple per-step recommendations with respect to a diverse set of future targets, we achieve notable offline and online benefits in performance, diversity, and efficiency.
  • Efficient Serving: We detail how we optimize autoregressive generation for optimal item recommendation, implement the serving infrastructure, and allow for horizontal scaling to serve hundreds of millions of users. To our knowledge, this study is the first to disclose the details of infrastructure optimization for industry-scale generative retrieval systems.
2. Related Work
2.1. Sequential Recommendation Systems
Sequential recommendation systems operate on a sequence of items with which a user has interacted (Kang and McAuley, 2018). Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018; Sun et al., 2019). Recent studies have harnessed the extensive knowledge of pre-trained language models (Hou et al., 2022; Li et al., 2023).
In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pancha et al., 2022). The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models). Due to its seamless integration with existing recommendation models, this approach has been widely adopted industry-wide (Hansen et al., 2020; Xu et al., 2022; Xia et al., 2023; Si et al., 2024).
Another approach in sequential recommendation systems is the generative approach, wherein the model learns to generate the items to recommend. Rajput et al. (2023) used transformers to generate items by generating a sequence of semantic ID tokens that are obtained by Residual Quantization with Variational Auto Encoder (RQ-VAE) (Lee et al., 2022; Singh et al., 2024). Zhai et al. (2024) augmented the sequence by adding heterogeneous items and also showed that the models can perform two tasks: ranking and retrieval. However, these studies have primarily concentrated on generating a single item rather than exploring the generation of sequences. In contrast, our research seeks to generate a sequence of multiple items to fully realize the potential of the generative approach. Our method is capable of recommending a diverse set of items that can drive various types of user actions. While similar approaches have been explored in academic settings (Liu et al., 2024), such strategies have not yet been widely adopted in industry.
2.2. Multi-Objective Generation
Given the multifaceted nature of user exploration and business needs on a platform like Pinterest, we explore generation techniques that consider multiple objectives. While reinforcement learning (RL) algorithms have demonstrated this capability in industry settings (Xin et al., 2020), research in language modeling indicates that RL often requires significant amounts of tuning for optimal generative capability (Ouyang et al., 2022). Recent research has introduced more streamlined and scalable techniques grounded in supervised learning. These approaches employ goal-conditioned methods (Ghosh et al., 2021) or utilize prompting strategies (Xin et al., 2022). Building on these, we develop an outcome-conditioned sequential recommendation system to serve billions of items and millions of users.
3. Problem Setup
In this section, we outline the problem setup for the task of a sequential recommendation system for a set of users $\mathcal{U}$, items $\mathcal{I}$, and actions $\mathcal{A}$. Our aim is to produce an ordered set of item recommendations $R(u,t)\subset\mathcal{I}$ of size $N$ for a user $u\in\mathcal{U}$ at a given timestep $t$. Each item $i$ has an item type $T(i)$, accounting for all kinds of items that the user can interact with (e.g., Pins). We define the set of all item types $\mathcal{T}$ as $\mathcal{T}=\{T(i)\mid i\in\mathcal{I}\}$.
Conditioned on the user $u$, we denote their historical user activity sequence ending at timestep $t_{\mathrm{max}}$ as $H(u,t_{\mathrm{max}})$, a chronologically-ordered heterogeneous sequence of items. For each item in the sequence $i_{t}\in H(u,t_{\mathrm{max}})$, the item $i_{t}$ is interacted by user $u$ at timestep $t\leq t_{\mathrm{max}}$ and its item type $T(i)\in\mathcal{T}$. We assume each interaction with an item $i_{t}$ corresponds to an action $a(i_{t})\in\mathcal{A}$.
To learn a sequential recommendation system, we propose a sequence model $f_{\theta}$ parameterized by $\theta$. To optimize the model parameters $\theta$, we leverage a set of features $\mathcal{F}(t)$ for each item type $t\in\mathcal{T}$. These may either be learned end-to-end with the sequence model or provided as learned or raw signals. The sequence model is consequently optimized using these features across the set of user sequences $H(u,t_{\mathrm{max}})$ for all $u\in\mathcal{U}$.
4. PinRec
In this section, we discuss the architecture, methodology, and serving strategies for the proposed recommender system, PinRec. Leveraging multi-token, autoregressive generation with outcome conditioning, we present an expressive, diverse, and performant transformer-based generative retrieval technique. We further discuss how PinRec is served in real-time through an in-house implementation of GPT-2 powered by NVIDIA Triton.
4.1. Representation Learning for Items
A critical prerequisite for effective optimization of a sequential recommendation system is an effective input and output representation for the heterogeneous set of items. We denote the representation of some item $i\in\mathcal{I}$ of type $T_{i}=T(i)$ as $f_{T_{i}}(i)$, where there exist representation models $f_{t}:\mathcal{F}(t)\to\mathcal{R}$ for each item type $t\in\mathcal{T}$ mapping the input feature space to the chosen representation space $\mathcal{R}$. For our purposes, we consider two item types ($\mathcal{T}$): Pins and search queries, which represent two core entities at Pinterest.
While some work has often leveraged ID-based representation spaces (Singh et al., 2024), recent work by Liu et al. (2024) and Yang et al. (2024) demonstrates that semantic IDs can lead to poor performance due to representational collapse. To compound onto this, given a larger set of item types $|\mathcal{T}|$, it may be challenging to optimize representations across many different types in a compatible representational space. On the contrary, Agarwal et al. (2024) observe that real-valued vector representations (i.e., $\mathcal{R}\subseteq\mathbb{R}^{d}$) generalize across item types.
[…]
As input to each embedder, we featurize the item $i$ given the chosen set of features $\mathcal{F}(T(i))$ and concatenate the features into a vector $\mathbf{v}_{i}$. To featurize search queries, we use the trained OmniSearchSage query representation (Agarwal et al., 2024), and to featurize Pins, we use the trained PinSage representation (Ying et al., 2018). Given that Pins are the core entity of Pinterest, we choose to implement a large, sharded set of hash-based embedding tables to store a learned ID embedding for each item. The output of the input embedder for item $i$ is a real-valued, $L_{2}$ normalized embedding $\mathbf{i}=f_{t}(\mathbf{v}_{i})$.
4.2. Sequential Modeling Architecture
To construct a sequential recommendation system, we leverage a causal transformer decoder architecture. Provided a sequence of input embeddings, we repeatedly apply multi-head attention followed by feedforward layers. Critically, we apply a causal attention mask at each token, i.e., preventing tokens at time $t$ from attending to any token at $t^{\prime}>t$. The output of the transformer $f_{\theta}(H(u,t_{\mathrm{max}}))$ is a sequence of hidden states from the last layer, and we apply an output head to obtain representations of recommended items.
We base our transformer decoder architecture on the GPT-2 model. We use an in-house implementation that optimizes existing solutions (e.g., HuggingFace [^1] https://huggingface.co/openai-community/gpt2), while allowing for additional functionality. For example, we incorporate the surface for each user interaction and the raw timestamp, following Pancha et al. (2022).
To generate the output representations for item recommendation, we construct an output head $O$ as a small MLP. While the output head is typically conditioned on the transformer’s hidden state, its input can incorporate other forms of conditioning. To that end, we denote any “auxiliary conditioning” as $\Phi(H(u,t_{\mathrm{max}}))$, and we explore this further in Section 4.3.1. We normalize the outputs so they are in the same representational space as the item representations $\mathbf{i}$. To summarize, the output representation $\hat{\mathbf{i}}_{u,t}$ at timestep $t$ for user $u$, applies the output head to the transformer’s last hidden state and any auxiliary conditioning (Equation 1).
(1) $\hat{\mathbf{i}}_{u,t}=O(f_{\theta}(H(u,t_{\mathrm{max}})),\Phi(H(u,t_{\mathrm{max}})))$
4.3. Optimizing Item Recommendations
In this section, we detail our approach to optimizing our item representations with a sequential objective. The standard choice to learn a distribution over personalized item recommendations is simply to employ a “next-token” prediction objective. Beyond directly taking into account the ground-truth user activity and engagement distribution, the next-token objective is effective, efficient, and ubiquitous. Following Pancha et al. (2022), we employ a generic sampled softmax loss to optimize our output item representations against our next item targets, using in-batch and random negatives, and with temperature scaling. Since in-batch negatives are sampled from a biased distribution based on our sequences, we employ sample probability correction through estimating the conditional frequency of each of the items $Q(\mathbf{i}_{c})=P(\mathbf{i}_{c}\mid\hat{\mathbf{i}}_{u,t})$ (Bengio and Senecal, 2008) with a count-min sketch (Cormode and Muthukrishnan, 2005).
Based on the vector representations of the output item representation $\hat{\mathbf{i}}_{u,t}$ and a candidate item $\mathbf{i}_{c}$, we define their similarity function $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{\mathbf{i}_{c}})$, applying a correction to account for sampling bias. The sampled softmax loss defined in Equation 2 maximizes the similarity of our output item representation $\hat{\mathbf{i}}_{u,t}$ and the next item representation $\mathbf{i}_{u,t+1}$, relative to in-batch and random negative examples $\mathbf{N}$.
(2) $L_{\mathrm{s}}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1})=-\log\left(\frac{\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1}))}{\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1}))+\sum_{\mathbf{i}_{n}\in\mathbf{N}}\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{n}))}\right)$
Our empirical next item loss function, over a user sequence $H(u,t_{\mathrm{max}})$ and sampled negatives $\mathbf{N}$, maximizes next item similarity across the entire sequence, as shown in Equation 3 (Pancha et al., 2022).
(3) $L_{\theta}^{\mathrm{next}}(H(u,t_{\mathrm{max}}))=\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t}L_{s}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1})$
To optimize the model parameters $\theta$ and consequently learn an optimal set of item recommendations, we minimize the empirical loss across user sequences and sampled negatives in our dataset $\mathcal{D}$.
(4) $\arg\min_{\theta}\mathbb{E}_{H(u,t_{\mathrm{max}})\sim\mathcal{D}}[L_{\theta}^{\mathrm{next}}(H(u,t_{\mathrm{max}}))]$
4.3.1. Outcome-Conditioned Generation
Although we optimize with respect to item targets conditioned on the user’s history, the next item loss provides no additional contextual information for the output representation by itself. For instance, given a user who has a propensity for low-effort engagement, the standard next token objective tends to generate recommendations that reinforce this same behavior. Though prior approaches leverage outcome-based loss weighting to resolve this, it requires domain knowledge and nevertheless requires significant amounts of tuning to achieve the desired outcome for each user. Alternatively, RL approaches are fraught with training instability and/or inefficiency and a significant need for hyperparameter tuning (Schulman et al., 2017). On the contrary, we propose reframing this problem by simply conditioning the generation of item recommendations directly on the desired outcomes (i.e., outcome or goal-conditioning).
To that end, we additionally condition the output head $O$ of the model on the desired outcome. To enable representing different outcomes, we employ a set of learnable embeddings for each possible action outcome (e.g., user repins, user clicks, etc.). By conditioning the output generation on a representation of the desired outcome (i.e., $\Phi(H(u,t_{\mathrm{max}}))$ represents $a_{\mathrm{desired}}\in\mathcal{A}$), we can generate outcome-specific, personalized item representations $\hat{\mathbf{i}}_{u,t}$.
At training time, we leverage user sequences to optimize the output representation with respect to the targets. Since we cannot dictate the desired outcome in offline data post-hoc, we optimize with respect to outcomes achieved in the data (Ghosh et al., 2021). Hence, we can define the auxiliary conditioning during training time as $\Phi(H(u,t_{\mathrm{max}}))=a(i_{t+1})$, corresponding to the target item $i_{t+1}$.
During inference, we can specify the desirability of each outcome and condition our generation according to these preferences. When the system is deployed, this allows us to induce these desirable outcomes from users. Formally, we define a per-action budget distribution $\mathcal{B}:\mathcal{A}\to\mathbb{R}$ that maps each action to a non-negative budget such that $\sum_{a\in\mathcal{A}}\mathcal{B}(a)=1$. Hence, as the budget for any action grows, it becomes more desirable. Based on the budget $\mathcal{B}$ and the desired numbers of recommended items $N$, we can either deterministically or stochastically allocate a fixed number of items for each action during generation (e.g., based on the expectation, each action $a$ can deterministically be allocated $\mathcal{B}(a)\cdot N$ items).
4.3.2. Windowed Multi-Token Generation
One of the principal assumptions in the next token objective is that there is strict ordering between the elements of the sequence. However, that reasoning is not necessarily applicable to social media platforms, where a user may engage with suggestions on their feed in an arbitrary order within a session. Moreover, a user may engage with certain items later on within a session (or perhaps at the very end) that they may have come across at the very beginning of the session. The next token objective simply ignores these crucial facets of recommendation systems for platforms such as Pinterest.
Motivated by the absence of strict ordering in sequential user engagement, we extend the next-token prediction objective. We assume that there exists a timestep window of size $\Delta$ within which user activity is roughly time-equivalent (i.e., their order within that window is not significant). To address the latter issue of delayed user engagement of an item relative to the current timestep, we employ an efficient multi-token objective which enables prediction beyond the next token at subsequent timesteps $t^{\prime}\geq t+1$.
Often, multi-token prediction for language models involves predicting a handful (e.g., 5-10) of successive tokens at timesteps $t+1$ to $t+k$ after the current token at timestep $t$ (Gloeckle et al., 2024). However, this leads to a similar issue of penalizing any predictions that do not match the target at exactly some timestep $t^{\prime}\in[t+1,t+k]$, even if a matching target may exist within the window $[t+1,t+\Delta]$. To incorporate the time-equivalent window size $\Delta$, we propose a windowed multi-token objective. That is, for any future timestep window of size $\Delta$ starting at some timestep $t^{\prime}>t$, all items that are engaged within that window are targets for our prediction task.
We choose to condition the output generation at timestep $t$ on both the target timestep $t^{\prime}>t$ and its action outcome (though it can work without outcome-conditioning). That is, we extend $\Phi(H(u,t_{\mathrm{max}}))$ to represent $\{t^{\prime},a(i_{t^{\prime}})\}$ at training time. Similarly to outcome-conditioning, we can decide $t^{\prime}$ (and the corresponding desired outcomes based on $\mathcal{B}$) at inference time as desired.
Our mathematical formulation for the windowed multi-token objective is shown in Equation 5, where $p_{t^{\prime}}$ denotes the chosen distribution over target timestep $t^{\prime}$ given $t$ (e.g., a uniform distribution from $[t+1,t_{\mathrm{lim}}]$ for some $t_{\mathrm{lim}}$). Note that a non-zero likelihood of sampling $t^{\prime}=t+1$ additionally enables this multi-token objective to perform next-token prediction and autoregressive generation.
(5) $\begin{split}L_{\theta}^{\mathrm{mt}}&(H(u,t_{\mathrm{max}}))=\\ &\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t}\mathbb{E}_{t^{\prime}\sim p_{t^{\prime}}(\cdot\mid t)}\Big{[}\min\{L_{s}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,i})\mid\forall i\in[t^{\prime},t^{\prime}+\Delta]\}\Big{]}\end{split}$
Intuitively, the optimization problem associated with the windowed multi-token objective can be posed as a question: can we sequentially predict any future items from the set of items that the user will engage with between timestep $t^{\prime}$ and $t^{\prime}+\Delta$? We find that this optimization is more intuitively aligned with the properties of user interaction on Pinterest compared to next-token prediction, while allowing for significantly reduced serving latency due to multiple items being predicted at a single token.
4.4. Serving
Figure 2. Serving flow for PinRec system (green boxes represent services, blue rectangles represent transmitted data, purple cylinders represent indexed data, and beige boxes are steps within the PinRec  Nvidia Triton ensemble).
PinRec is implemented with an in-house, optimized implementation of GPT-2, and served with NVidia Triton Inference Server’s ensemble setting and Python backend, to perform autoregressive generative retrieval. This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with learned ID embeddings, autoregressive inference for item generation, and Faiss retrieval (Figure 2).
4.4.1. Signal Fetching
Our Signal Service uses a Lambda architecture, which combines both batch and real-time processing. The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, and stores the sequences and features for each user ID. The real-time component is backed by a RocksDB key-value store (Prasad, 2024) that allows us to efficiently fetch user engagements since the end time of the batch signal. The batch and real-time signals are de-duplicated and returned to PinRec. On the search surface, model inference also includes the search query issued by the user at request time as the final element in the sequence. To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision. These are later de-quantized to FP16 for inference.
4.4.2. Learned ID Embeddings
As discussed in Section 4.1, our pin embedder includes an ID embedding table. Effectively, it is represented as a large in-memory table that maps each pin ID’s hashes to an embedding, followed by some MLP layers. As this ID embedder takes up significant space in memory, especially when replicated, it is separated out from the main PinRec model and runs in its own CPU memory-optimized service. It is served with torchscript to limit latencies to single-digit milliseconds.
4.4.3. Autoregressive Item Generation
To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer with a KV Cache. Depending upon whether the model is outcome-conditioned or multi-token, we are able to sample or generate multiple representations respectively. After autoregressive generation, we use Faiss to retrieve a set of recommended items of size $N$ corresponding to the generated representations. Several optimizations are required for efficient inference:
CUDA Graphs.
To limit overhead from kernel launches and include other in-built PyTorch optimizations, we use CUDA Graphs with torch.compile to maximally compile the main components of the model.
KV Cache.
During the prefill stage of the transformer, the $K$ and $V$ tensors for the user’s known sequence are cached in a per-batch-request KV Cache. On subsequent decode steps, these are reused, and only new sequence elements are processed to produce $K$ and $V$ for the additional sequence dimension.
Compression.
Given that we generate multiple output representations with autoregressive generation, we observe there are often embeddings that are similar to one another. During Faiss retrieval, these are likely to collapse into a similar set of output items. To optimize this, we “compress” each generated embedding into a prior uncompressed embedding, in the order it is generated, if it is more similar than a given cosine similarity threshold. Compression significantly helps reduce the median number of embeddings generated by each model, with an average ratio of 6:1 for an unconditioned PinRec model and a smaller ratio of 3:1 for the multi-token, outcome-conditioned variant.
Since we wish to fetch $N$ total items, we allocate these proportionally to each compressed embedding based on the total number of embeddings that were compressed into it. In the case of outcome-conditioning, we must perform this allocation independently for each action, in adherence with the chosen budget $\mathcal{B}$. The set of compressed embeddings alongside the allocations are passed to the Faiss retrieval stage.
4.4.4. Faiss Retrieval
(a) Recent user history specified to produce results in Subfigure 3(b).
(b) Results for the user sequence specified in Subfigure 3(a).
Figure 3. Item recommendations from PinRec-OC for a user history with cat-related search queries and pin interactions. History in reverse chronological order. Additional examples in Section A.1.
Budget is: {repin: 0.4, outbound_click: 0.1, grid_click: 0.4, long_grid_click: 0.05, other: 0.05}. Generation runs for 6 steps (1 prefill, 5 decode).
Our nearest-neighbor retrieval involves a Faiss IVF-HNSW index running on CPU hosts, with inner product as the similarity metric. As with the main model, this is also served with NVidia Triton with batching enabled. We run $k$NN retrieval to fetch the allocated number of Pins for each embedding, with an overfetch factor for later de-duplication. Sample results are visualized for a manually specified user history in Figure 3.
5. Offline Evaluation
In this section, we dissect the performance of the proposed sequential recommendation system, PinRec,  on generative retrieval across a range of offline evaluations. Our experimental setting includes interactions on three major surfaces for the Pinterest platform, namely Homefeed, Related Pins, and Search. We demonstrate that outcome-conditioned generation allows for improved controllability and that multi-token generation leads to significantly improved diversity, performance, and efficiency.
5.1. Offline Evaluation Metrics
We construct the evaluation set using interactions for a set of users disjoint from the training set and evaluate whether the proposed technique can produce a representation to retrieve future item embeddings among a set of 1M random items (“index”). In such tasks, recall @ $k$ is a standard performance metric, defined as the proportion of target items correctly retrieved by the model within the top-$k$ ranked items across an index. However, standard recall assumes the existence of a single target per sample. As mentioned in Section 4.3.2, such an approach fails to capture the assumption of a lack of strict ordering between user activity within a small range of interactions $\Delta$. Further, when the goal of the retrieval model is candidate generation for a downstream ranking model, the task naturally has more than one unique plausible target, making standard recall less relevant.
We address these shortcomings by evaluating model performance on an unordered generalization of recall @ $k$. Given the autoregressively predicted item representations across all users $\mathbf{\hat{I}}$, target item embeddings across all users $\mathbf{I}$, and random item embeddings $\mathbf{N}$. Differently from standard recall, for each user $u\in\mathcal{U}$, unordered recall compares a set of predicted $\mathbf{\hat{I}}_{u}$ with a set of targets $\mathbf{I}_{u}$ of fixed size $\Delta$, calculating the proportion of targets $\mathbf{I}_{u}$ where there exists some $\mathbf{\hat{I}}_{u}$ that is within the top-$k$ ranked items. We formalize the notion of unordered recall in Equation 6.
(6) $\begin{split}\mathrm{ur}_{k}(\mathbf{\hat{I}},\mathbf{I},\mathbf{N})\coloneq\frac{1}{|\mathbf{\mathbf{\hat{I}}}|}\times\sum_{i=1}^{|\mathbf{\hat{I}}|}\mathbf{1}\left\{\min_{j,z\in[\Delta]}\left|\{\mathbf{n}\in\mathbf{N}\mid\mathbf{\hat{I}}_{ij}^{\top}\mathbf{n}\geq\mathbf{\hat{I}}_{ij}^{\top}\mathbf{I}_{iz})\}\right|\leq k\right\}\end{split}$
Note that since we perform autoregressive generation without teacher forcing, the transformer is not conditioned on any of the targets at any point.
Beyond unordered retrieval performance, we examine the diversity of the recommendations. After retrieving the top-$k$ predicted items for each user, we compute the proportion of unique items retrieved across the evaluation set. This allows us to quantify the degree to which learned representations are biased towards popular or frequently occurring Pins in the dataset.
5.2. Offline Results
In this section, we compare PinRec with an existing sequence-based user modeling representation, Pinnerformer (Pancha et al., 2022), in terms of offline evaluation metrics. Further, we quantify the impacts of our design on the performance and diversity of recommendations.
5.2.1. Comparison with Prior Method
To validate that PinRec outperforms existing sequential techniques in terms of item recommendation in offline evaluation, we compare its performance to PinnerFormer. We evaluate three different variants of PinRec.
  • PinRec-UC: We do not condition on anything (i.e., $\Phi=\emptyset$) and do not employ multi-token prediction.
  • PinRec-OC: We condition on the target action (“outcome conditioned”) and do not employ multi-token prediction.
  • PinRec-{MT, OC}: We condition on the target action and employ multi-token prediction.
To ensure that the comparison between PinRec-{MT, OC} and other methods is fair, we adjust the number of autoregressive generation steps for multi-token generation to ensure that the total number of generated embeddings is held constant. As in training (where we cannot modify the evaluation dataset post-hoc), we condition on the target item’s action for outcome-conditioned variants.
Our comparison in Table 1 shows that PinRec outperforms a similarly sized Pinnerformer across the board. Importantly, outcome-conditioning demonstrates notable improvements over PinRec-UC (unconditioned) in unordered recall @ 10 (e.g., +2-3% on Homefeed and Related Pins). Above all, PinRec-{MT, OC} shows improvements over every method, with over +10% lift on Homefeed, over +20% on Related Pins, and +30% on search over PinRec-UC.
Table 1. Comparison of PinnerFormer and PinRec variants (unordered recall @ 10) across major surfaces at Pinterest.
Model Homefeed Related Pins Search
PinnerFormer 0.461 0.412 0.257
PinRec-UC 0.608 0.521 0.350
PinRec-OC 0.625 0.537 0.352
PinRec-{MT, OC} 0.676 0.631 0.450
5.2.2. Outcome-Conditioned Generation
We seek to understand whether we can achieve “controllability” between the desired action and true action; for instance, does conditioning on a “repin” generate an output representation that truly aligns with content that will be repinned? If we are able to control the outcome through our item recommendations, then we will observe that for the specific action that we condition on, there are significant gains in offline recall. Conversely, we expect that the recall for items corresponding to other actions will decrease since our action-conditioned generation should ideally induce only that chosen action. As a baseline, we compare to the unconditioned variant of PinRec, which does not take into account the desired action.
For simplicity, we allocate the entire budget entirely to one action and perform offline evaluation conditioning across each action. We show the result of this analysis in Figure 4. Whenever we condition on an action outcome, we show notable gains over an unconditioned baseline for that action. Specifically, matching outcome conditioning yields 1.9% recall lift for true repins, 1.4% on true grid clicks, 4.8% on long grid clicks, and 6.2% on outbound clicks.
On the other hand, we observe decreases in recall compared to an unconditioned baseline if the true action does not correspond to the desired action, indicating that our conditioning achieves the desired result of producing outcome-specific representations. Especially for outbound clicks, we observe that conditioning on any other actions leads to massive drops in unordered recall.
Figure 4. Percentage lift in unordered recall for PinRec-OC over PinRec-UC when conditioning on the desired action, stratified by the actual action taken by the user.
5.2.3. Utility of Multi-Token Generation
Beyond the improvement in performance, we evaluate the benefits of multi-token generation in terms of efficiency, diversity, and performance of generations relative to the number of chosen generations per timestep. We examine the tradeoff between generating more representations per timestep versus longer autoregressive generation, while keeping the number of output embeddings constant. Since the strongest offline performance aside from the multi-token variant is by PinRec-OC, we consider it as a baseline. For simplicity, we average the unordered recall @ 10 across the three surfaces.
We show the tradeoff between performance, diversity, and efficiency with the number of embeddings we generate in each generation step using multi-token generation in Figure 5. While next-token generation (i.e., 1 embedding per step) is the most inefficient and yields the poorest performance (where we perform similarly to PinRec-OC), we observe significant benefits across the board in increasing the number of embeddings per step. At 16 generations per step, we demonstrate $\sim$10x reduction in latency over PinRec-OC, with improved unordered recall (+16.0%) and diversity (+21.3%).
Figure 5. Tradeoff between unordered recall @ 10 (top), proportion unique (bottom), and generation latency (right) with the number of multi-token generations per step.
6. Online A/B Experiments
We perform online A/B experiments to evaluate PinRec as a candidate generator, measuring its impact on user engagement and confirming its online controllability. The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other candidate generators, to produce the final results shown to users. Since we employ a heterogeneous sequence of Pins and search queries, we focus on the Homefeed and Search surfaces. On these surfaces, we provide personalized item recommendations through PinRec, leading to significant lifts in action metrics, session metrics, and intent fulfillment.
6.1. Homefeed
6.1.1. Experiment Configuration
In the Homefeed online experiment, we run five groups exposed to global traffic for PinRec-OC, with varying action budgets $\mathcal{B}$. The budgets are fixed for outbound click (0.1), long grid clicks (0.05), and other unspecified actions (0.05). The remaining budget is split between the repin and grid click engagements, in steps of 0.2 ranging from 0 to 0.8. We similarly configure five groups for PinRec-{OC, MT}, with the same budgets. We include control groups without PinRec.
6.1.2. Metrics
Pinterest uses “session fulfillment” as an informative and holistic whole-funnel metric to capture (a) the upper funnel of user intent expression (e.g., when a user clicks on a pin in the grid), (b) the middle funnel of curating Pins (e.g., repin), and (c) the lower funnel of realizing intent (e.g., clicking on a product pin to visit its website). Any of the aforementioned actions results in a session being marked as “fulfilled”, whereas the absence of these results in an “unfulfilled” session.
6.1.3. Results
In Figure 6, we visualize a Pareto front of the repin rate (relative to impressions) and grid click rate for the PinRec candidate generator by itself, with absolute numbers hidden due to non-public information. By reducing the grid click budget and correspondingly increasing the repin budget, we see the expected result as the grid click rate drops and repin rate increases (i.e., we demonstrate online “controllability”). Similarly, we observe a significant correlation between the budgets and the overall A/B experiment metric lifts for Homefeed grid clicks and repins in Figure 7. Further, we note that across both variants in Figure 7, for $\mathcal{B}(\text{repin})\geq 0.6$, there are statistically significant overall online lifts in Homefeed repin. We achieve statistically significant lifts in Homefeed grid click for all budget groups on each variant.
(a) PinRec-OC
(b) PinRec-{MT, OC}
Figure 6. Visualization of Pareto fronts for online rate of grid click versus repin for items recommended only by PinRec variants, with error bars denoted in grey.
(a) PinRec-OC
(b) PinRec-{MT, OC}
Figure 7. Visualization of Pareto fronts for online lift % in Homefeed grid clicks versus repin for PinRec variants.
To examine the holistic impact of our candidate generators, we examine the lift in fulfilled sessions for each PinRec variant, selecting the group which maximizes this metric. To demonstrate notable site-wide, session-level, and global metrics improvements, we present the CUPED-adjusted  (Deng et al., 2013) percentage lift in metrics across Homefeed and the entire site (Table 2) for the groups that maximize fulfilled sessions. Relative to the control group, the unconditioned variant of PinRec (PinRec-UC) demonstrates improvements in unfulfilled sessions and in Homefeed and site-wide grid clicks. In addition to those lifts, the outcome-conditioned variant (PinRec-OC) shows significant lift in site-wide fulfilled sessions. The multi-token variant, PinRec-{MT, OC}, significantly increased successful site sessions, time spent on site, and site-wide grid clicks, which we believe represent the most comprehensive site-wide improvements.
Table 2. CUPED-adjusted metrics improvements from online A/B experiments for PinRec variants as candidate generators in Homefeed.
Type Metric UC OC {MT, OC}
Overall Fulfilled Sessions +0.02% +0.21% +0.28%
Metrics Time Spent -0.02% +0.16% +0.55%
Unfulfilled Sessions -0.28% -0.89% -0.78%
Action Site-wide Grid Clicks +0.58% +1.76% +1.73%
Metrics Homefeed Grid Clicks +1.87% +4.01% +3.33%
6.2. Search
On the Search surface, we evaluate the unconditioned variant of PinRec. As shown in Table 3, PinRec-UC improves search metrics significantly relative to a control group. We observe session-wide search “fulfillment” increase, i.e., when a search session results in positive interaction such as repins or long outbound clicks.
Table 3. CUPED-adjusted metrics improvements from online A/B experiments for PinRec as candidate generator in Search retrieval.
Metric Type Metric PinRec-UC
Overall Metrics Search Fulfillment Rate +0.48%
Action Metrics Search Repins +0.84%
Search Grid Clicks +1.46%
7. Conclusion
We introduced PinRec, a generative retrieval system developed for Pinterest’s Homefeed and Search feed. To the best of our knowledge, this paper represents the first comprehensive study on deploying generative retrieval in a web-scale application. The incorporation of two key innovations — outcome-conditioned generation and multi-token inference — enabled PinRec to enhance several business metrics without substantial cost increases.
Several promising future directions can be pursued. First, applying generative approaches to ranking, as explored by Zhai et al. (Zhai et al., 2024), could be beneficial. Second, optimizing input sequences to the model, akin to prompt engineering in large language models, is another avenue worth exploring. Third, developing language model alignment techniques similar to Ouyang et al. (Ouyang et al., 2022) could enhance the use of engagement signals from users. Overall, this paper highlights a promising avenue for utilizing generative retrieval in large-scale recommender systems.
Acknowledgements.
We thank collaborators such as Deepak Agarwal, Divyansh Agarwal, Edoardo Botta, Bee-Chung Chen, Bowen Deng, Pong Eksombatchai, Kurchi Subhra Hazra, Se Won Jang, Saurabh Vishwas Joshi, Krishna Kamath, Sujay Khandagale, James Li, Cristian Lopez, Matt Madrigal, Lei Pan, and Ruijia Wang for their critical contributions to our work and for facilitating productive discussions.
Appendix A Appendix
A.1. Visualization of Multi-Token Impact
In this section, we present some additional qualitative examples through visualizations of the retrieved examples for some sample user histories across the windowed multi-token and non-multi-token variants. Through these visualizations, we are able to demonstrate a reasonable and consistent qualitative difference in the types of pins retrieved, justifying the improved session-wide and sitewide online metrics for PinRec-{MT, OC}.
Given a user history, we display two examples of retrieved pins in Figure 8 and Figure 9 for PinRec-{MT, OC} and PinRec-{OC}. Specifically, note that PinRec-{MT, OC} captures less-recent interactions for users compared to PinRec-OC. This is expected as the multi-token model predicts over a longer future time window and is less biased toward recency. PinRec-{MT, OC} can therefore produce more diverse results capturing more medium-term interests. On the other hand, since the PinRec-{OC} is naturally biased towards the short term, it fails to capture potential longer term trends that may induce engagement.
Note that these trends reflect the same takeaways from offline and online evaluation. Offline, we observe an improvement in diversity as the number of multi-token predictions is increased, which is consistently above the non-multi-token variant of PinRec. Online, we see that the “feed” presented by PinRec-{MT, OC} is of higher quality due to its diversity and visual appeal/aesthetic. These are critical qualities for a platform like Pinterest.
(a) Recent user history manually specified to produce results in Subfigures LABEL:subfig:nails_tattoos_user_non_mt and LABEL:subfig:nails_tattoos_user_mt. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
(b) Results for PinRec-OC candidate generator for user history in subfigure above. Notably only includes Pins related to nails, the user’s most recent interaction.
(c) Results for PinRec-{MT, OC} candidate generator for user history in subfigure above. Includes a mix of Pins related to both tattoos and nails.
Figure 8. Comparison of results for a manually-specified user history.
(a) Recent user history manually specified to produce results in Subfigures 9(b) and 9(c). In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
(b) Results for PinRec-OC candidate generator for user history in Subfigure 9(a). Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
(c) Results for PinRec-{MT, OC} candidate generator for user history in Subfigure 9(a). Includes a mix of Pins related to nurseries and food.
Figure 9. Comparison of results for another manually-specified user history.

### v2

@@
 Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (Vaswani et al., 2017), and then use those tokens to retrieve items.
-Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018; Li et al., 2023; Yang et al., 2024).
+Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018a; Li et al., 2023; Yang et al., 2024).
 Nevertheless, to our knowledge, generative retrieval has not been attempted as successfully for large-scale industrial applications yet.
@@
 To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges.
-Most methods developed within academic settings (Kang and McAuley, 2018; Hou et al., 2022; Li et al., 2023) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic.
+Most methods developed within academic settings (Kang and McAuley, 2018a; Hou et al., 2022; Li et al., 2023) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic.
 While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (Zhai et al., 2024) and do not demonstrate optimal optimization for different target outcomes.
@@
 2.1. Sequential Recommendation Systems
-Sequential recommendation systems operate on a sequence of items with which a user has interacted (Kang and McAuley, 2018).
-Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018; Sun et al., 2019).
+Sequential recommendation systems operate on a sequence of items with which a user has interacted (Kang and McAuley, 2018a).
+Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018a; Sun et al., 2019).
 Recent studies have harnessed the extensive knowledge of pre-trained language models (Hou et al., 2022; Li et al., 2023).
@@
 5.2. Offline Results
-In this section, we compare PinRec with an existing sequence-based user modeling representation, Pinnerformer (Pancha et al., 2022), in terms of offline evaluation metrics.
+In this section, we compare PinRec with an existing sequence-based user modeling representations, Pinnerformer (Pancha et al., 2022), SASRec (Kang and McAuley, 2018b), and TIGER (Rajput et al., 2023), in terms of offline evaluation metrics.
 Further, we quantify the impacts of our design on the performance and diversity of recommendations.
@@
 Above all, PinRec-{MT, OC} shows improvements over every method, with over +10% lift on Homefeed, over +20% on Related Pins, and +30% on search over PinRec-UC.
-Table 1. Comparison of PinnerFormer and PinRec variants (unordered recall @ 10) across major surfaces at Pinterest.
+Table 1. Comparison of baselines and PinRec variants (unordered recall @ 10) across major surfaces at Pinterest.
 Model Homefeed Related Pins Search
-PinnerFormer 0.461 0.412 0.257
+TIGER (Rajput et al., 2023) 0.208 0.230 0.090
+SASRec (Kang and McAuley, 2018b) 0.382 0.426 0.142
+PinnerFormer (Pancha et al., 2022) 0.461 0.412 0.257
 PinRec-UC 0.608 0.521 0.350
@@
 Appendix A Appendix
-A.1. Visualization of Multi-Token Impact
+A.1. Additional Analysis
+In this section, we perform additional experiments to quantitatively and qualitatively show that the design decisions made for PinRec are justified and that it performs across the state-of-the-art architectures.
+Additionally, we provide more online analyses of the performance across different user cohorts and tail items.
+A.1.1. HSTU Architecture
+Though PinRec relies primarily on the well-known transformer architecture, we note that our technique is completely architecture-agnostic.
+Specifically, outcome conditioning and multi-token prediction can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements.
+To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (Zhai et al., 2024), as opposed to the multi-head attention-based causal transformer.
+Note that we do not leverage the ranking task or any sequence sampling explored in (Zhai et al., 2024) as it is not relevant for our specific use case.
+For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
+The results are shown in Table 4, which clearly shows that HSTU is roughly 3% worse on HomeFeed (effectively negating the improvement with respect to the unconditioned transformer-based PinRec) and competitive on the related pins surface.
+However, there is a substantial drop in performance on the search surface, which is a critical component of the user experience at Pinterest.
+We believe that this may be related to specific architectural elements of HSTU.
+Table 4. Comparison of HSTU and transformer architectures for PinRec, where PinRec-OC leverages the transformer as presented in the main text.
+Model Homefeed Related Pins Search
+PinRec-OC 0.625 0.537 0.352
+HSTU (OC) 0.596 0.539 0.179
+Though we observe a sharp drop in the search surface, HSTU performs roughly on the same order of magnitude as the traditional transformer architecture, demonstrating that our ideas perform across different architectures.
+This validates that our ideas are robust to architectures across the board.
+A.1.2. Ablating Negative Sampling
+Following from Pancha et al.
+(2022), we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function.
+Though we ultimately end up with similar conclusions as Pancha et al.
+(2022), we show these findings for maximum clarity regarding our design choices.
+Random Negatives
+Though in-batch negatives are a critical component of our negative corpus, we construct a random negative corpus consisting of a random set of possible items.
+Given that our retrieval problem is ultimately across a random selection of pins, we wish for PinRec to show retrieval capability across this corpus.
+We perform several ablations with random negatives, with in-batch negatives constant at a non-zero value.
+For these analyses, we focus on Homefeed (the most critical surface) and present results aggregated across all surfaces as well.
+We show results in Table 5, where there are moderate gains in unordered recall if there is a non-zero number of in-batch negatives present.
+Specifically, on the most critical surface of Homefeed, we see notable improvement over the baseline in terms of unordered recall.
+Table 5. Percentage lift in unordered recall using 16K random negatives, relative to using only 8K random negatives.
+Surface % Lift in Recall @ 10
+Homefeed +2.5%
+All +1.3%
+To summarize, while random negatives are certainly an important piece of the puzzle, they cannot substitute in-batch negatives.
+When used in conjunction, scaling the number of negatives yields notable improvement, especially on Homefeed.
+In-Batch Negatives
+As shown in the previous subsection, disabling in-batch negatives entirely leads to significantly poorer metrics, showing their importance.
+Can we improve performance offline by scaling the batch size and consequently, the number of in-batch negatives?
+To examine this, we fix the number of random negatives and ablate the number of in-batch negatives through the batch size.
+Table 6. Percentage lift in unordered recall @ 10 through scaling the number of in-batch negatives (relative to baseline of 1x) with random negatives fixed.
+# In-Batch Negs. % Lift (HF) % Lift (all)
+1.5x +4.2% +3.3%
+2x +8.4% +5.4%
+12x +18.3% +12.1%
+We show results in Table 6, which demonstrates a clear trend that shows that a greater number of in-batch negatives and batch size leads to better performance, especially for the Homefeed surface.
+These are similar in directional improvements to random negatives, but we observe greater improvements with in-batch negatives, which are likely more relevant as negative examples.
+A.1.3. Ablating Parameter Scaling
+One of the recent explorations within recommendation systems has been scaling to improve performance, in an attempt to replicate the success of these efforts for language modeling.
+To explore this hypothesis, we scale the number of parameters within PinRec across two fronts: the “vocabulary” and the model itself.
+While model scaling is clear and involves scaling up the embedding size, the number of transformer layers, and so on, scaling the vocabulary for an embedding-based generative retrieval technique involves improving the input representations.
+Specifically, as mentioned in the main text, we implement “ID” embeddings for pins, that effectively serve as additional representation for the “pin vocabulary”.
+Model Parameters
+As presented in the main text, PinRec (all variants) have roughly 100 million transformer parameters, corresponding to the base form or size of GPT-2.
+To ablate the effect of the size of the transformer, we attempt to follow the different scales employed by GPT-2 (base, medium, large, XL) as they have shown success in terms of the factors by which the embedding size, number of layers, and number of heads have scaled.
+Note that due to computational constraints, we reduce the scaling in terms of the number of layers, which increases GPU memory footprint significantly.
+Note that we present lift in offline unordered recall @ 10 for the outcome conditioned variant of PinRec below, but we observe similar trends for the unconditioned model as well.
+Table 7. Percentage lift in unordered recall through expanding the transformer at various GPT-2 scales relative to base ( 100 million parameters).
+Model % Lift in Recall @ 10
+Medium -0.8%
+Large -0.9%
+XL +2.2%
+Though we observe some gains in Table 7 from scaling the number of transformer parameters to around 1 billion (XL-sized), these become intractable to serve in real time, requiring offline batch inference or significantly higher cost and more optimization.
+For the remaining sizes, we observe insignificant lifts (note that less than 1% in lift magnitude can be often explained through noise).
+Hence, we opt to not use larger sized transformers.
+Vocabulary Parameters
+As presented in the main text, the ID embeddings consist of roughly 10 billion parameters.
+These are a core component of the model since they provide additional representational power for the most important entity type at Pinterest: Pins.
+To ablate the importance of these ID embeddings, we examine the performance in offline recall across all surfaces with and without ID embeddings.
+For this analysis, we focus on Homefeed (the most critical surface) and present results aggregated across all surfaces as well in Table 8.
+Table 8. Percentage lift through adding ID embeddings (10 billion parameters) over no ID embedding-based vocabulary.
+Surface % Lift in Recall @ 10
+Homefeed +14.0%
+All +14.4%
+Evidently, the improvement in recall is substantial and indicates that the scaling of vocabulary to learn an embedding for each unique ID (roughly, as it is hash-based) rather than only content-based features is beneficial.
+A.1.4. Tail User and Item Analysis
+One of the potential drawbacks of outcome conditioning would be for “tail” or “fresh” users, i.e., those which have lesser activity or less propensity for high-effort engagement.
+Critically, our method intends to optimize for outcomes across all cohorts of users, including such users.
+In the same vein, we wish to optimize for “tail” items, which may receive a lesser number of impressions, in an effort to distribute content to improve corpus coverage.
+User Cohort Grid Clicks Outbound CTs
+UC OC UC OC
+New +2.60% +2.62% +3.98% +5.78%
+Resurrected +1.56% +1.86% +1.98% +1.72%
+(a) Online lift (%) for grid clicks and outbound clickthroughers for new and resurrected users.
+Metric UC OC
+Corpus Coverage (unique pins) +6.2% +17.3%
+1-Imp Pin Rec. (tail items) +5.8% +7.7%
+(b) Relative lift (vs. 2-tower model) for tail item and corpus coverage metrics.
+Table 9. Online experimental results for different user cohorts and tail item metrics, across PinRec-UC and PinRec-OC.
+In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 9(a).
+Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users.
+In fact, across new users, OC performs better on outbound CTs.
+To examine performance for tail items, we compare to a production 2-tower model in an online setting in Table 9.
+On corpus coverage, adding PinRec-UC/OC surfaces 6.2%/17.3% more unique pins relative to the 2-tower.
+For long-tail items with 1-impression, PinRec-UC/OC achieve 5.8%/7.7% absolute lifts respectively, relative to the aforementioned 2-tower model.
+Evidently, across the board, PinRec surfaces more of the corpus and surfaces more tail items compared to control, with outcome conditioning showing even more significant lifts over control in both regards.
+We believe the evidence suggests that our generative retrieval solution is equal to or better across inactive/new users and tail items relative to existing control, and perhaps more importantly, it shows that outcome conditioning can help in boosting performance in these criteria, which is one of its goals.
+A.1.5. Serving Efficiency
+One of the critical aspects of our proposed generative retrieval technique is an effort to maintain serving efficiency.
+With our custom infrastructure built from off-the-shelf serving technologies such as NVidia Triton Inference Server and optimization techniques like torch.compile, we achieve a system that can be served in real-time.
+In this section, we delve deeper into the memory and time requirements, comparing our method with prior methods deployed in production.
+On the memory front, all variants of PinRec (unconditioned, outcome-conditioned with or without multi-token generation) use about 1.6GiB of GPU memory including KVCache and CUDA graphs, without any significant memory usage increase for multi-token prediction or outcome conditioning.
+Excluding the KVCache and CUDA graphs, we use roughly the same amount of memory (2̃50 MiB) relative to other techniques (i.e., two-tower models).
+In terms of latency, at 80 QPS, the online and offline latency for PinRec-OC/OC, MT are p50 40ms/p90 65ms (6 steps).
+On the other hand, across only 3 steps for PinRec-UC, we observe a latency of p50 23ms and p90 46ms (note that this produces significantly less embeddings and less compressed embeddings than either of the above techniques, though it is lower latency).
+Though the latency is roughly 3-4x higher (p50) than a traditional two-tower solution, we note that parallelizing it with other RPCs results in ¡1% added end-to-end increase in latency.
+A.2. Visualization of Multi-Token Impact
 In this section, we present some additional qualitative examples through visualizations of the retrieved examples for some sample user histories across the windowed multi-token and non-multi-token variants.
@@
 These are critical qualities for a platform like Pinterest.
-(a) Recent user history manually specified to produce results in Subfigures LABEL:subfig:nails_tattoos_user_non_mt and LABEL:subfig:nails_tattoos_user_mt.
-In reverse chronological order.
-Note a mix of nail and tattoo-related queries and engagements.
+(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
 (b) Results for PinRec-OC candidate generator for user history in subfigure above. Notably only includes Pins related to nails, the user’s most recent interaction.
@@
 Figure 8. Comparison of results for a manually-specified user history.
-(a) Recent user history manually specified to produce results in Subfigures 9(b) and 9(c). In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
-(b) Results for PinRec-OC candidate generator for user history in Subfigure 9(a). Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
-(c) Results for PinRec-{MT, OC} candidate generator for user history in Subfigure 9(a). Includes a mix of Pins related to nurseries and food.
+(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
+(b) Results for PinRec-OC candidate generator for given user history. Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
+(c) Results for PinRec-{MT, OC} candidate generator for given user history. Includes a mix of Pins related to nurseries and food.
 Figure 9. Comparison of results for another manually-specified user history.

### v4

@@
 Additionally, PinRec incorporates multi-token generation to enhance output diversity while optimizing generation.
-Our experiments demonstrate that PinRec can successfully balance performance, diversity, and efficiency, delivering a significant positive impact to users using generative models.
+Our experiments demonstrate that PinRec can successfully balance performance, diversity, and efficiency, delivering a significant positive impact such as +2% sitewide clicks and +4% search repins.
 This paper marks a significant milestone in generative retrieval, as it presents, to our knowledge, the first rigorous study on implementing generative retrieval at the scale of Pinterest.
@@
 1. Introduction
-Pinterest is a visual discovery platform where over 550 million monthly active users search for, save, and shop for ideas, represented by Pins (Pinterest, 2024a).
-Given the massive volume of Pins — for instance, users save over 1.5 billion Pins per week (Pinterest, 2024b) — it is essential to provide personalized recommendations in Pinterest feeds and search results.
+Pinterest is a visual discovery platform where over 550 million monthly active users search for, save, and shop for ideas, represented by Pins (pins).
+Given the massive volume of Pins — for instance, users save over 1.5 billion Pins per week (pins2) — it is essential to provide personalized recommendations in Pinterest feeds and search results.
 To achieve this, recommender systems often leverage a two-stage approach: retrieval and ranking.
@@
 As an alternative to two-tower models, recent work has proposed generative retrieval.
-Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (Vaswani et al., 2017), and then use those tokens to retrieve items.
-Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018a; Li et al., 2023; Yang et al., 2024).
-Nevertheless, to our knowledge, generative retrieval has not been attempted as successfully for large-scale industrial applications yet.
+Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (vaswani2017attention), and then use those tokens to retrieve items.
+Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (KangM18; Recformer23; UnifyingGenRetrieval24).
 Using generative retrieval models for large scale recommender systems poses many challenges at the scale of Pinterest.
-Given their increased complexity and sequential generation, generating a set of recommended items requires higher computational cost than computing an embedding with two-tower models.
-Further, the generative model should be capable of efficiently optimizing for different engagement outcomes because (a) business needs are constantly evolving to optimize combinations of different metrics and (b) as importantly, we wish for users to engage with content in the most optimal way for their growth on the platform.
-For instance, we would want to approach recommendations for a user with a propensity for low-effort engagement differently than for core users; through which recommendations could we enable the maximum progress in their user journey on Pinterest?
-To answer this question requires a malleable technique that can adapt to different circumstances, goals, and demands.
-Finally, the generated item sequences must possess sufficient diversity to facilitate users’ exploration of varied ideas; one-step generation of a single item representation may not suffice.
+Firstly, the generative model should be capable of efficiently optimizing for different engagement outcomes such as maximizing social actions like saves and comments versus passive engagements like clicks, because business needs are constantly evolving to optimize combinations of different metrics.
+Secondly, the generated item sequences must possess sufficient diversity to facilitate users’ exploration of varied ideas; one-step generation of a single item representation may not suffice.
+Thirdly, given their increased complexity and sequential generation, generating a set of recommended items requires higher computational cost than computing an embedding with two-tower models.
 To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges.
-Most methods developed within academic settings (Kang and McAuley, 2018a; Hou et al., 2022; Li et al., 2023) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic.
-While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (Zhai et al., 2024) and do not demonstrate optimal optimization for different target outcomes.
-In this paper, we present PinRec, a generative retrieval technique for real-world recommender systems.
-Unlike previous methods, PinRec addresses and resolves three crucial criteria for generation of optimal recommendations: multifaceted generative capabilities, computational efficiency, and item diversity.
+Most methods developed within academic settings (KangM18; UniS4Rec22; Recformer23) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic.
+While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (HSTU24; Tiger23) and are not capable of optimizing multiple target outcomes.
+Moreover, these methods do not address the challenge of generating a sequence of diverse items.
+Because generating multiple items sequentially is computationally expensive, existing approaches typically generate only the next item rather than a full set at once.
+Figure 2. Modeling diagram of PinRec
+In this paper, we present PinRec, a generative retrieval technique for real-world recommender systems (Figure 2).
+PinRec builds on other generative retrieval models, leveraging transformers to perform next token prediction (KangM18; HSTU24; Tiger23).
+Unlike previous methods, however, PinRec addresses and resolves three aforementioned challenges for generation of optimal recommendations: multi-objective generative capabilities, computational efficiency, and item diversity.
 Our main contributions in the paper are summarized below.
-  • Outcome-Conditioned Generative Retrieval: Given evolving business needs and to allow for different types of user exploration, we propose multifaceted generative capabilities through outcome-conditioned generation.
-Across offline and online evaluation, we demonstrate that customizing item recommendations to the desired outcome allows for designers to provide recommendations that successfully result in those specified outcomes.
-  • Windowed, Multi-Token Prediction: We propose an efficient multi-token prediction objective tailored to the structure of the sequential recommendation task, where prior next-token objectives have often lacked representational power.
-Instead, by optimizing multiple per-step recommendations with respect to a diverse set of future targets, we achieve notable offline and online benefits in performance, diversity, and efficiency.
+  • Outcome-Conditioned Generative Retrieval: PinRec employs outcome-conditioned generation, enabling the optimization of multiple business metrics.
+For each outcome (e.g., business metrics such as user saves or clicks), PinRec generates the next item token that is most likely to maximize that specific outcome, and then retrieves candidate items to balance these objectives according to the desired trade-offs (Fig.
+2).
+Through both offline and online evaluations, we show that tailoring item recommendations to particular outcomes enables designers to effectively achieve those specified objectives.
+  • Multi-Token Generation for Efficient Sequence Generation: One obstacle to generating an entire sequence of items for retrieval is that traditional models typically produce only one item at a time.
+To address this, we propose a multi-token generation framework (Multitoken24), in which PinRec generates multiple future items in parallel, as illustrated in Figure 2.
+This approach significantly reduces the number of steps needed to produce diverse output embeddings, resulting in substantial offline and online improvements in performance, diversity, and efficiency.
   • Efficient Serving: We detail how we optimize autoregressive generation for optimal item recommendation, implement the serving infrastructure, and allow for horizontal scaling to serve hundreds of millions of users.
 To our knowledge, this study is the first to disclose the details of infrastructure optimization for industry-scale generative retrieval systems.
+We will make the model code publicly available upon publication.
 2. Related Work
-2.1. Sequential Recommendation Systems
-Sequential recommendation systems operate on a sequence of items with which a user has interacted (Kang and McAuley, 2018a).
-Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018a; Sun et al., 2019).
-Recent studies have harnessed the extensive knowledge of pre-trained language models (Hou et al., 2022; Li et al., 2023).
-In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pancha et al., 2022).
-The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models).
-Due to its seamless integration with existing recommendation models, this approach has been widely adopted industry-wide (Hansen et al., 2020; Xu et al., 2022; Xia et al., 2023; Si et al., 2024).
-Another approach in sequential recommendation systems is the generative approach, wherein the model learns to generate the items to recommend.
-Rajput et al.
-(2023) used transformers to generate items by generating a sequence of semantic ID tokens that are obtained by Residual Quantization with Variational Auto Encoder (RQ-VAE) (Lee et al., 2022; Singh et al., 2024).
-Zhai et al.
-(2024) augmented the sequence by adding heterogeneous items and also showed that the models can perform two tasks: ranking and retrieval.
-However, these studies have primarily concentrated on generating a single item rather than exploring the generation of sequences.
-In contrast, our research seeks to generate a sequence of multiple items to fully realize the potential of the generative approach.
-Our method is capable of recommending a diverse set of items that can drive various types of user actions.
-While similar approaches have been explored in academic settings (Liu et al., 2024), such strategies have not yet been widely adopted in industry.
-2.2. Multi-Objective Generation
-Given the multifaceted nature of user exploration and business needs on a platform like Pinterest, we explore generation techniques that consider multiple objectives.
-While reinforcement learning (RL) algorithms have demonstrated this capability in industry settings (Xin et al., 2020), research in language modeling indicates that RL often requires significant amounts of tuning for optimal generative capability (Ouyang et al., 2022).
-Recent research has introduced more streamlined and scalable techniques grounded in supervised learning.
-These approaches employ goal-conditioned methods (Ghosh et al., 2021) or utilize prompting strategies (Xin et al., 2022).
-Building on these, we develop an outcome-conditioned sequential recommendation system to serve billions of items and millions of users.
+Sequential Recommendation Systems: Sequential recommendation systems operate on a sequences of items with which a user has interacted (KangM18).
+Earlier work in academia has used recurrent neural networks (RNN) (GRU4REC16), but later work has often adopted transformers (KangM18; BERT4REC19).
+Recent studies have harnessed the extensive knowledge of pre-trained language models (UniS4Rec22; Recformer23).
+In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pinnerformer22).
+The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models) (Spotify20; rethink; Transact23; TWINV224).
+Another approach in sequential recommendation systems is the generative approach, where the model learns to generate the items to recommend.
+Tiger23 used transformers to generate items by producing a sequence of semantic ID tokens (RQVAE22; SemanticIDGeneral24).
+HSTU24 also proposed a generative retrieval method.
+However, these studies have mainly focused on generating a single item to optimize a single action.
+In contrast, our research aims to generate sequences of multiple items to optimize for multiple outcomes.
+Multi-Objective Generation: We discuss generative models that consider multiple objectives.
+While RL algorithms have shown this capability in industry (SelfsupervisedRL20), studies in language modeling suggest RL often needs extensive tuning for optimal generation (PPO22).
+Recent research has introduced more streamlined and scalable techniques based on supervised learning, such as goal-conditioned methods (IterativeSupervised21) and prompting strategies (PromptingRL22).
+PinRec builds on these approaches to enhance scalability.
 3. Problem Setup
-In this section, we outline the problem setup for the task of a sequential recommendation system for a set of users $\mathcal{U}$, items $\mathcal{I}$, and actions $\mathcal{A}$.
-Our aim is to produce an ordered set of item recommendations $R(u,t)\subset\mathcal{I}$ of size $N$ for a user $u\in\mathcal{U}$ at a given timestep $t$.
-Each item $i$ has an item type $T(i)$, accounting for all kinds of items that the user can interact with (e.g., Pins).
-We define the set of all item types $\mathcal{T}$ as $\mathcal{T}=\{T(i)\mid i\in\mathcal{I}\}$.
-Conditioned on the user $u$, we denote their historical user activity sequence ending at timestep $t_{\mathrm{max}}$ as $H(u,t_{\mathrm{max}})$, a chronologically-ordered heterogeneous sequence of items.
-For each item in the sequence $i_{t}\in H(u,t_{\mathrm{max}})$, the item $i_{t}$ is interacted by user $u$ at timestep $t\leq t_{\mathrm{max}}$ and its item type $T(i)\in\mathcal{T}$.
-We assume each interaction with an item $i_{t}$ corresponds to an action $a(i_{t})\in\mathcal{A}$.
-To learn a sequential recommendation system, we propose a sequence model $f_{\theta}$ parameterized by $\theta$.
-To optimize the model parameters $\theta$, we leverage a set of features $\mathcal{F}(t)$ for each item type $t\in\mathcal{T}$.
-These may either be learned end-to-end with the sequence model or provided as learned or raw signals.
-The sequence model is consequently optimized using these features across the set of user sequences $H(u,t_{\mathrm{max}})$ for all $u\in\mathcal{U}$.
+We formulate the generative sequential recommendation task as follows.
+Consider a system with users $\mathcal{U}$, items $\mathcal{I}$, actions $\mathcal{A}$, and surfaces $\mathcal{S}$.
+Our objective is to generate an ordered list of $N$ item recommendations $R(u,t)\subset\mathcal{I}$ for user $u\in\mathcal{U}$ at timestep $t$.
+Item Types and Actions.
+Each item $i\in\mathcal{I}$ belongs to an item type $T(i)\in\mathcal{T}$, where $\mathcal{T}=\{T(i)\mid i\in\mathcal{I}\}$ represents the set of all item types in the system (e.g., Pins, boards, search queries).
+Users can perform various actions $\mathcal{A}=\{\text{click},\text{save},\text{share},\text{comment},\ldots\}$ on these items across different surfaces.
+Each interaction occurs on a specific surface $s\in\mathcal{S}$, where surfaces represent different recommendation products, such as homefeed, search results, board pages, and related pin feed.
+User History.
+For each user $u$, we maintain their interaction history $H(u,t_{\text{max}})=\langle(i_{1},a_{1},s_{1},t_{1}),(i_{2},a_{2},s_{2},t_{2}),\ldots,(i_{m},a_{m},s_{m},t_{m})\rangle$, where each tuple $(i_{j},a_{j},s_{j},t_{j})$ represents an interaction with item $i_{j}\in\mathcal{I}$ via action $a_{j}\in\mathcal{A}$ on surface $s_{j}\in\mathcal{S}$ at timestep $t_{j}\leq t_{\text{max}}$.
+The sequence is chronologically ordered such that $t_{1}\leq t_{2}\leq\ldots\leq t_{m}$.
+Sequential Recommendation Model.
+We propose a parameterized sequence model $f_{\theta}:\mathcal{H}\times\mathcal{T}\times\mathcal{S}\rightarrow\mathbb{R}^{|\mathcal{I}|}$ that maps a user’s interaction history, item types, and surface context to item scores, where $\mathcal{H}$ denotes the space of all possible interaction histories.
+The model parameters $\theta$ are optimized over the collection of user interaction sequences $\{H(u,t_{\text{max}})\mid u\in\mathcal{U}\}$ to maximize the likelihood of observed user interactions while accounting for the heterogeneous nature of actions and surfaces.
+This enables the model to learn nuanced behavioral patterns that depend on both what users interact with and how and where they interact with it.
 4. PinRec
-In this section, we discuss the architecture, methodology, and serving strategies for the proposed recommender system, PinRec.
-Leveraging multi-token, autoregressive generation with outcome conditioning, we present an expressive, diverse, and performant transformer-based generative retrieval technique.
-We further discuss how PinRec is served in real-time through an in-house implementation of GPT-2 powered by NVIDIA Triton.
-4.1. Representation Learning for Items
-A critical prerequisite for effective optimization of a sequential recommendation system is an effective input and output representation for the heterogeneous set of items.
-We denote the representation of some item $i\in\mathcal{I}$ of type $T_{i}=T(i)$ as $f_{T_{i}}(i)$, where there exist representation models $f_{t}:\mathcal{F}(t)\to\mathcal{R}$ for each item type $t\in\mathcal{T}$ mapping the input feature space to the chosen representation space $\mathcal{R}$.
-For our purposes, we consider two item types ($\mathcal{T}$): Pins and search queries, which represent two core entities at Pinterest.
-While some work has often leveraged ID-based representation spaces (Singh et al., 2024), recent work by Liu et al.
-(2024) and Yang et al.
-(2024) demonstrates that semantic IDs can lead to poor performance due to representational collapse.
-To compound onto this, given a larger set of item types $|\mathcal{T}|$, it may be challenging to optimize representations across many different types in a compatible representational space.
-On the contrary, Agarwal et al.
-(2024) observe that real-valued vector representations (i.e., $\mathcal{R}\subseteq\mathbb{R}^{d}$) generalize across item types.
-[…]
-As input to each embedder, we featurize the item $i$ given the chosen set of features $\mathcal{F}(T(i))$ and concatenate the features into a vector $\mathbf{v}_{i}$.
-To featurize search queries, we use the trained OmniSearchSage query representation (Agarwal et al., 2024), and to featurize Pins, we use the trained PinSage representation (Ying et al., 2018).
-Given that Pins are the core entity of Pinterest, we choose to implement a large, sharded set of hash-based embedding tables to store a learned ID embedding for each item.
-The output of the input embedder for item $i$ is a real-valued, $L_{2}$ normalized embedding $\mathbf{i}=f_{t}(\mathbf{v}_{i})$.
-4.2. Sequential Modeling Architecture
-To construct a sequential recommendation system, we leverage a causal transformer decoder architecture.
-Provided a sequence of input embeddings, we repeatedly apply multi-head attention followed by feedforward layers.
-Critically, we apply a causal attention mask at each token, i.e., preventing tokens at time $t$ from attending to any token at $t^{\prime}>t$.
-The output of the transformer $f_{\theta}(H(u,t_{\mathrm{max}}))$ is a sequence of hidden states from the last layer, and we apply an output head to obtain representations of recommended items.
-We base our transformer decoder architecture on the GPT-2 model.
-We use an in-house implementation that optimizes existing solutions (e.g., HuggingFace [^1] https://huggingface.co/openai-community/gpt2), while allowing for additional functionality.
-For example, we incorporate the surface for each user interaction and the raw timestamp, following Pancha et al.
-(2022).
-To generate the output representations for item recommendation, we construct an output head $O$ as a small MLP.
-While the output head is typically conditioned on the transformer’s hidden state, its input can incorporate other forms of conditioning.
-To that end, we denote any “auxiliary conditioning” as $\Phi(H(u,t_{\mathrm{max}}))$, and we explore this further in Section 4.3.1.
-We normalize the outputs so they are in the same representational space as the item representations $\mathbf{i}$.
-To summarize, the output representation $\hat{\mathbf{i}}_{u,t}$ at timestep $t$ for user $u$, applies the output head to the transformer’s last hidden state and any auxiliary conditioning (Equation 1).
-(1) $\hat{\mathbf{i}}_{u,t}=O(f_{\theta}(H(u,t_{\mathrm{max}})),\Phi(H(u,t_{\mathrm{max}})))$
-4.3. Optimizing Item Recommendations
-In this section, we detail our approach to optimizing our item representations with a sequential objective.
-The standard choice to learn a distribution over personalized item recommendations is simply to employ a “next-token” prediction objective.
-Beyond directly taking into account the ground-truth user activity and engagement distribution, the next-token objective is effective, efficient, and ubiquitous.
-Following Pancha et al.
-(2022), we employ a generic sampled softmax loss to optimize our output item representations against our next item targets, using in-batch and random negatives, and with temperature scaling.
-Since in-batch negatives are sampled from a biased distribution based on our sequences, we employ sample probability correction through estimating the conditional frequency of each of the items $Q(\mathbf{i}_{c})=P(\mathbf{i}_{c}\mid\hat{\mathbf{i}}_{u,t})$ (Bengio and Senecal, 2008) with a count-min sketch (Cormode and Muthukrishnan, 2005).
-Based on the vector representations of the output item representation $\hat{\mathbf{i}}_{u,t}$ and a candidate item $\mathbf{i}_{c}$, we define their similarity function $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{\mathbf{i}_{c}})$, applying a correction to account for sampling bias.
-The sampled softmax loss defined in Equation 2 maximizes the similarity of our output item representation $\hat{\mathbf{i}}_{u,t}$ and the next item representation $\mathbf{i}_{u,t+1}$, relative to in-batch and random negative examples $\mathbf{N}$.
-(2) $L_{\mathrm{s}}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1})=-\log\left(\frac{\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1}))}{\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1}))+\sum_{\mathbf{i}_{n}\in\mathbf{N}}\exp(s(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{n}))}\right)$
-Our empirical next item loss function, over a user sequence $H(u,t_{\mathrm{max}})$ and sampled negatives $\mathbf{N}$, maximizes next item similarity across the entire sequence, as shown in Equation 3 (Pancha et al., 2022).
-(3) $L_{\theta}^{\mathrm{next}}(H(u,t_{\mathrm{max}}))=\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t}L_{s}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,t+1})$
-To optimize the model parameters $\theta$ and consequently learn an optimal set of item recommendations, we minimize the empirical loss across user sequences and sampled negatives in our dataset $\mathcal{D}$.
-(4) $\arg\min_{\theta}\mathbb{E}_{H(u,t_{\mathrm{max}})\sim\mathcal{D}}[L_{\theta}^{\mathrm{next}}(H(u,t_{\mathrm{max}}))]$
-4.3.1. Outcome-Conditioned Generation
-Although we optimize with respect to item targets conditioned on the user’s history, the next item loss provides no additional contextual information for the output representation by itself.
-For instance, given a user who has a propensity for low-effort engagement, the standard next token objective tends to generate recommendations that reinforce this same behavior.
-Though prior approaches leverage outcome-based loss weighting to resolve this, it requires domain knowledge and nevertheless requires significant amounts of tuning to achieve the desired outcome for each user.
-Alternatively, RL approaches are fraught with training instability and/or inefficiency and a significant need for hyperparameter tuning (Schulman et al., 2017).
-On the contrary, we propose reframing this problem by simply conditioning the generation of item recommendations directly on the desired outcomes (i.e., outcome or goal-conditioning).
-To that end, we additionally condition the output head $O$ of the model on the desired outcome.
-To enable representing different outcomes, we employ a set of learnable embeddings for each possible action outcome (e.g., user repins, user clicks, etc.).
-By conditioning the output generation on a representation of the desired outcome (i.e., $\Phi(H(u,t_{\mathrm{max}}))$ represents $a_{\mathrm{desired}}\in\mathcal{A}$), we can generate outcome-specific, personalized item representations $\hat{\mathbf{i}}_{u,t}$.
-At training time, we leverage user sequences to optimize the output representation with respect to the targets.
-Since we cannot dictate the desired outcome in offline data post-hoc, we optimize with respect to outcomes achieved in the data (Ghosh et al., 2021).
-Hence, we can define the auxiliary conditioning during training time as $\Phi(H(u,t_{\mathrm{max}}))=a(i_{t+1})$, corresponding to the target item $i_{t+1}$.
-During inference, we can specify the desirability of each outcome and condition our generation according to these preferences.
-When the system is deployed, this allows us to induce these desirable outcomes from users.
-Formally, we define a per-action budget distribution $\mathcal{B}:\mathcal{A}\to\mathbb{R}$ that maps each action to a non-negative budget such that $\sum_{a\in\mathcal{A}}\mathcal{B}(a)=1$.
-Hence, as the budget for any action grows, it becomes more desirable.
-Based on the budget $\mathcal{B}$ and the desired numbers of recommended items $N$, we can either deterministically or stochastically allocate a fixed number of items for each action during generation (e.g., based on the expectation, each action $a$ can deterministically be allocated $\mathcal{B}(a)\cdot N$ items).
-4.3.2. Windowed Multi-Token Generation
-One of the principal assumptions in the next token objective is that there is strict ordering between the elements of the sequence.
-However, that reasoning is not necessarily applicable to social media platforms, where a user may engage with suggestions on their feed in an arbitrary order within a session.
-Moreover, a user may engage with certain items later on within a session (or perhaps at the very end) that they may have come across at the very beginning of the session.
-The next token objective simply ignores these crucial facets of recommendation systems for platforms such as Pinterest.
-Motivated by the absence of strict ordering in sequential user engagement, we extend the next-token prediction objective.
-We assume that there exists a timestep window of size $\Delta$ within which user activity is roughly time-equivalent (i.e., their order within that window is not significant).
-To address the latter issue of delayed user engagement of an item relative to the current timestep, we employ an efficient multi-token objective which enables prediction beyond the next token at subsequent timesteps $t^{\prime}\geq t+1$.
-Often, multi-token prediction for language models involves predicting a handful (e.g., 5-10) of successive tokens at timesteps $t+1$ to $t+k$ after the current token at timestep $t$ (Gloeckle et al., 2024).
-However, this leads to a similar issue of penalizing any predictions that do not match the target at exactly some timestep $t^{\prime}\in[t+1,t+k]$, even if a matching target may exist within the window $[t+1,t+\Delta]$.
-To incorporate the time-equivalent window size $\Delta$, we propose a windowed multi-token objective.
-That is, for any future timestep window of size $\Delta$ starting at some timestep $t^{\prime}>t$, all items that are engaged within that window are targets for our prediction task.
-We choose to condition the output generation at timestep $t$ on both the target timestep $t^{\prime}>t$ and its action outcome (though it can work without outcome-conditioning).
-That is, we extend $\Phi(H(u,t_{\mathrm{max}}))$ to represent $\{t^{\prime},a(i_{t^{\prime}})\}$ at training time.
-Similarly to outcome-conditioning, we can decide $t^{\prime}$ (and the corresponding desired outcomes based on $\mathcal{B}$) at inference time as desired.
-Our mathematical formulation for the windowed multi-token objective is shown in Equation 5, where $p_{t^{\prime}}$ denotes the chosen distribution over target timestep $t^{\prime}$ given $t$ (e.g., a uniform distribution from $[t+1,t_{\mathrm{lim}}]$ for some $t_{\mathrm{lim}}$).
-Note that a non-zero likelihood of sampling $t^{\prime}=t+1$ additionally enables this multi-token objective to perform next-token prediction and autoregressive generation.
-(5) $\begin{split}L_{\theta}^{\mathrm{mt}}&(H(u,t_{\mathrm{max}}))=\\ &\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t}\mathbb{E}_{t^{\prime}\sim p_{t^{\prime}}(\cdot\mid t)}\Big{[}\min\{L_{s}(\mathbf{\hat{i}}_{u,t},\mathbf{i}_{u,i})\mid\forall i\in[t^{\prime},t^{\prime}+\Delta]\}\Big{]}\end{split}$
-Intuitively, the optimization problem associated with the windowed multi-token objective can be posed as a question: can we sequentially predict any future items from the set of items that the user will engage with between timestep $t^{\prime}$ and $t^{\prime}+\Delta$?
-We find that this optimization is more intuitively aligned with the properties of user interaction on Pinterest compared to next-token prediction, while allowing for significantly reduced serving latency due to multiple items being predicted at a single token.
+In this section, we describe the model architecture and serving infrastructure for PinRec.
+We start with an overview of the model, including outcome-conditioned generation and multi-token prediction.
+Next, we explain how items in the sequence are encoded and processed, and how the training loss is computed.
+Finally, we describe inference mode operation and provide details about our real-time serving architecture, which is powered by Nvidia Triton.
+4.1. Model Architecture
+We use a transformer decoder architecture as the backbone of our PinRec model.
+The user interaction history is first featurized and fed into a stack of transformer layers along with learned position embeddings to capture temporal ordering.
+At each layer, we apply causal attention masking to ensure that tokens at time $t$ can only attend to tokens at positions $t^{\prime}\leq t$, preventing information leakage from future interactions.
+The stack of $L$ transformer layers outputs a sequence of contextualized hidden states: $\mathbf{h}_{u,t}=\text{TransformerStack}^{(L)}(\mathbf{x}_{u,1:t})$.
+where $\mathbf{x}_{u,1:t}$ represents the featurized user interaction sequence up to timestep $t$, and $\mathbf{h}_{u,t}$ is the resulting hidden state at position $t$ from the final layer.
+These hidden states are then processed through an output head $O$ to generate the next item predictions.
+Outcome-Conditioned Generation.
+Unlike language modeling, where the output sequence receives a single reward value (PPO22), users interact with recommended items through diverse engagement types—saving, sharing, commenting, or clicking.
+To effectively optimize across these varied engagement metrics, generative retrieval methods must account for this behavioral diversity (MultiBehaviorSASRec24).
+Furthermore, practical deployment often requires specifying the target surface or context for the next recommendation, even when the user’s historical sequence spans multiple surfaces.
+To address these requirements, we design our model to accept conditioning signals at each generation step, enabling dynamic control over model behavior during inference.
+Specifically, we condition the output head $O$ on a desired set of outcomes, allowing the model to generate targeted recommendations.
+We implement this through learnable embeddings for each possible outcome type (e.g., engagement actions, surface constraints).
+By conditioning the generation process on these outcome representations, our model produces outcome-specific, personalized item representations:
+(1) $\hat{\mathbf{i}}_{u,t}=O(\mathbf{h}_{u,t},\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k})$
+where $\mathbf{h}_{u,t}$ is the hidden state at timestep $t$, $\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k}$ are the conditioning embeddings for different outcome types (e.g., $\mathbf{c}_{1}=\mathbf{e}_{\text{action}}$, $\mathbf{c}_{2}=\mathbf{e}_{\text{surface}}$), and $\hat{\mathbf{i}}_{u,t}$ is the resulting conditioned item representation.
+Temporal Multi-token Prediction.
+Multi-token prediction in language models involves predicting multiple successive tokens at positions $t+1$ to $t+k$ given the current position $t$ (gloeckle2024better).
+This is typically implemented using separate output heads for each future position.
+However, in recommender systems, previous research (Pinnerformer22) suggests that the difference in time between actions is much more important than the position index alone, because a single position increment can correspond to very different time intervals for active users versus dormant users.
+To address this, we propose a time-based multi-token approach, in which tokens are generated for each time interval $\delta$ (e.g., generating recommendations for the user for the present, 30 seconds later, 1 minute later, and so on).
+Specifically, we condition our output head on the relative temporal distance from the current action, enabling a single parameterized head to generate predictions at various future time intervals:
+(2) $\hat{\mathbf{i}}_{u,t+\delta}=O(\mathbf{h}_{u,t},\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k},\mathbf{e}_{\delta})$
+where $\mathbf{e}_{\delta}$ is the learnable embedding for temporal offset $\delta\in\{1,2,\ldots,k\}$, and $\hat{\mathbf{i}}_{u,t+\delta}$ represents the predicted item at timestep $t+\delta$.
+4.2. Input Sequence Representation
+Given the heterogeneous nature of user interactions $(i_{j},a_{j},s_{j},t_{j})$ in our interaction history $H(u,t_{\text{max}})$, we need to transform each element into appropriate representations for the transformer architecture.
+A crucial design decision is defining representations for heterogeneous item types.
+There are two main ways to represent items in a sequence: (1) by using dense embedding vectors (HSTU24; KangM18), and (2) by learning a sequence of discrete IDs for each item, known as Semantic IDs (Tiger23; SemanticIDGeneral24).
+We explored both approaches and ultimately chose to use dense embedding vectors.
+In our experiments, we observed that Semantic IDs frequently suffered from representational collapse—many items were assigned the same Semantic ID, as also noted in (MultiBehaviorSASRec24; Hourglass24)—and consistently produced lower evaluation metrics (see Table 1(UnifyingGenRetrieval24).
+While optimizing performance with Semantic IDs remains an interesting direction for future work, our core contributions are representation-agnostic.
+We now detail how each component of the interaction tuple $(i_{j},a_{j},s_{j},t_{j})$ is processed and integrated into our model architecture.
+Item Embeddings.
+We generate dense embeddings for different entity types ($\tau\in\mathcal{T}$) using type-specific embedders that are learned end-to-end with the model.
+Each item type $\tau$ has a dedicated embedder $f_{\tau}$ that maps type-specific features to the shared representation space.
+While we describe embedders for Pins and search queries in this work, the approach can be extended to other entity types such as boards, users, or topics.
+Pin Embeddings.
+For Pins, we combine semantic embedding features with learned Id embeddings (10.1145/3640457.3688053).
+The semantic representation uses the pre-trained OmniSage embedding (badrinath2025omnisage), which captures visual and textual content along with the engagement via the Pin-board graph.
+Additionally, we implement a large-scale ID embedding system where each Pin identifier is hashed into $k$ different hash functions, each mapping to embedding tables of size $d/k$.
+These $k$ embeddings are concatenated to form a $d$-dimensional ID embedding, providing a distributed representation that scales efficiently.
+The ID embeddings constitute sparse parameters that allow us to scale model capacity with minimal impact on serving latency and computational cost, as only a small subset of embeddings are accessed for any given batch of items.
+The concatenated OmniSage and ID embeddings are passed through 2-3 fully connected layers with ReLU activation and layer normalization, followed by $L_{2}$ normalization to produce the final Pin embedding $\mathbf{i}_{\text{pin}}$.
+Query Embeddings.
+For search queries, we use only semantic representations as ID embeddings have not provided value in offline metrics for this entity type.
+We leverage the pre-trained OmniSearchSage query representation (agarwal2024omnisearchsage) which captures semantic query understanding.
+This representation is passed through a similar MLP architecture (2-3 fully connected layers with ReLU activation and layer normalization) followed by $L_{2}$ normalization to generate the final query embedding $\mathbf{i}_{\text{query}}$.
+Temporal Embeddings.
+For each interaction timestep $t_{j}$, we extract absolute timestamp and time since previous action, capturing both global temporal context and local session dynamics.
+Our encoder applies sinusoidal transformations at multiple frequencies.
+For absolute timestamps, we use predefined periods $P=\{p_{1},p_{2},\ldots,p_{k}\}$ to capture daily, weekly, and seasonal patterns: $\sin(t_{\text{abs}}/p_{i}),\cos(t_{\text{abs}}/p_{i})$.
+For relative time differences, we use log-scale frequencies augmented with learnable phase parameters.
+Sequence Construction.
+The final input sequence to the transformer is constructed by concatenating item and temporal embeddings, then adding surface embeddings for each interaction.
+This creates a rich contextual representation that captures what users interacted with (item), when they interacted (temporal), and where the interaction occurred (surface) within the platform.
+4.3. Training Objective
+During training, we employ teacher forcing with a next-item prediction objective.
+The hidden states $\mathbf{h}_{u,t}$ are processed through the conditioned output head $O$ to generate item representations that predict future items in the sequence.
+For outcome-conditioned generation, we condition the output head on the actual values of the target item during training.
+For multi-token prediction, we sample $k$ positions from the next $N$ positions in the sequence to serve as prediction targets.
+Following Pinnerformer22, we compute similarity between the predicted item representation $\hat{\mathbf{i}}_{u,t}$ and candidate item $\mathbf{i}_{c}$ as $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{i}_{c})$, where $Q(\mathbf{i}_{c})$ is a bias correction term estimated using count-min sketches (cormode2005improved) to account for the non-uniform sampling of in-batch negatives.
+The sampled softmax loss maximizes similarity to target items relative to negative samples $\mathbf{N}$:
+(3) $L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}})=-\log\frac{\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}}))}{\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}}))+\sum_{\mathbf{i}_{n}\in\mathbf{N}}\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{n}))}$
+We average this loss over all positions in each user sequence. For standard next-token prediction, the target is simply the immediate next item:
+(4) $L_{\theta}^{\text{next}}(H(u,t_{\max}))=\frac{1}{|H(u,t_{\max})|}\sum_{t=1}^{t_{\max}-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
+However, this strict sequential assumption may not align well with user interaction patterns.
+User interaction sequences differ fundamentally from language: while they exhibit global structure reflecting long-term interests, they lack the local structural constraints found in natural language.
+On social media platforms, users engage with feed suggestions in arbitrary order within sessions, making strict positional dependencies less meaningful than in natural language modeling.
+To better accommodate these interaction patterns, we introduce an intra-feed relaxation that considers all items engaged within the same feed session as valid targets.
+Given a feed session $F_{s}=\{i_{t_{1}},i_{t_{2}},\ldots,i_{t_{n}}\}$ where $t_{1}<t_{2}<\ldots<t_{n}$, we define the relaxed target set for position $t_{j}$ as:
+$T_{t_{j}}^{\text{feed}}=\{i_{t_{k}}\mid t_{k}\in F_{s},t_{k}>t_{j}\}$
+This enables a relaxed objective that optimizes for predicting any future item within the same session:
+(5) $L_{\theta}^{\text{feed}}(H(u,t_{\mathrm{max}}))=\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t_{j}}\min_{i\in T_{t_{j}}^{\text{feed}}}L_{s}(\hat{\mathbf{i}}_{u,t_{j}},i)$
+where we minimize the loss to the best-matching item among all future items within the same feed session, rather than strictly requiring prediction of the immediate next item.
+Intuitively, this optimization can be posed as: can we predict any of the future items that the user will engage with within the same feed session?
+We find that this approach is more naturally aligned with user interaction patterns on Pinterest compared to strict next-token prediction.
+Finally, we minimize the expected loss across all user sequences in dataset $\mathcal{D}$:
+(6) $\arg\min_{\theta}\mathbb{E}_{H(u,t_{\max})\sim\mathcal{D}}[L_{\theta}^{\text{feed}}(H(u,t_{\max}))].$
+4.4. Inference
+Our PinRec model serves as a candidate generator within the broader recommendation stack, producing diverse item embeddings.
+The model operates in an autoregressive fashion to generate multiple candidate embeddings, which are then used to retrieve actual items through approximate nearest neighbor search.
+To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer. The specific generation process varies based on the model configuration:
+Unconditional Generation.
+For the baseline model, generation is straightforward: at each step, the model produces a single item embedding, which is appended to the sequence along with the target surface embedding and a predefined temporal offset (e.g., 10 seconds).
+This updated sequence is then used to generate the next embedding.
+Outcome-Conditioned Generation.
+For models with outcome conditioning, for each outcome we generate a candidate embedding, producing multiple candidate embeddings per generation step.
+For autoregressive continuation, we sample a value for each condition according to the requested distribution over each condition $D_{1},D_{2},\ldots,D_{n}$ and use the corresponding embedding as input for the next step (Fig.
+2).
+Temporal Multi-Token Generation.
+When using temporal multi-token prediction, we sample time deltas up to a specified maximum and generate multiple embeddings at different temporal offsets, combined with any specified outcome conditions.
+For the next autoregressive input, we use the embedding corresponding to the immediate next token.
+Across $n$ rollouts, this process generates varying numbers of embeddings depending on the model configuration and input settings.
+The generated embeddings are then used for candidate retrieval through Approximate Nearest Neighbor (ANN) search.
+Budget Allocation.
+To retrieve $N$ total items, we first allocate retrieval budget across all generated embeddings.
+For unconditional models, the budget is distributed equally among all embeddings.
+For outcome-conditioned models, allocation is performed independently for each action type according to the specified budget constraints $\mathcal{B}$, ensuring that each condition receives its designated proportion of the total retrieval budget.
+Embedding Compression.
+Given that autoregressive generation produces multiple output representations, we observe that many embeddings are often similar to one another.
+During ANN retrieval, these similar embeddings would collapse into overlapping sets of retrieved items, reducing diversity.
+To address this, we implement an embedding compression strategy: each generated embedding is merged into a previously uncompressed embedding (in generation order) if their cosine similarity exceeds a predefined threshold.
+When embeddings are compressed together, their allocated budgets are summed to determine the total retrieval budget for the compressed group.
+This candidate generation process enables PinRec to serve as a flexible first-stage retrieval system within Pinterest’s recommendation infrastructure.
+The generated candidates can be further refined by downstream ranking models, allowing the system to benefit from both the generative model’s ability to capture complex user patterns and the precision of specialized ranking algorithms.
+The outcome-conditioning and multi-token capabilities make the system particularly suitable for multi-objective optimization scenarios common in production recommendation systems.
 4.5. Serving
 Figure 3.
-Serving flow for PinRec system (green boxes represent services, blue rectangles represent transmitted data, purple cylinders represent indexed data, and beige boxes are steps within the PinRec  Nvidia Triton ensemble).
-PinRec is implemented with an in-house, optimized implementation of GPT-2, and served with NVidia Triton Inference Server’s ensemble setting and Python backend, to perform autoregressive generative retrieval.
+Serving flow for the PinRec system.
+Green boxes represent services, blue rectangles represent transmitted data, purple cylinders represent indexed data, and beige boxes are steps within the PinRec NVIDIA Triton ensemble.
+PinRec is implemented with an in-house, optimized implementation of transformer and served with NVIDIA Triton Inference Server’s ensemble setting and Python backend to perform autoregressive generative retrieval.
 This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with learned ID embeddings, autoregressive inference for item generation, and Faiss retrieval (Figure 3).
 4.5.1. Signal Fetching
-Our Signal Service uses a Lambda architecture, which combines both batch and real-time processing.
-The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, and stores the sequences and features for each user ID.
-The real-time component is backed by a RocksDB key-value store (Prasad, 2024) that allows us to efficiently fetch user engagements since the end time of the batch signal.
-The batch and real-time signals are de-duplicated and returned to PinRec.
-On the search surface, model inference also includes the search query issued by the user at request time as the final element in the sequence.
-To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision.
-These are later de-quantized to FP16 for inference.
+Our Signal Service uses a Lambda architecture that combines both batch and real-time processing.
+The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, storing the sequences and features for each user ID.
+The real-time component is backed by a RocksDB key-value store (kvstore) that efficiently retrieves user engagements occurring after the batch signal’s cutoff time.
+The batch and real-time signals are deduplicated and returned to PinRec.
+On surfaces with context (e.g., Search, Related Pins), model inference also includes the context (e.g., search query, closeup Pin) issued by the user at request time as the final element in the sequence.
+To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision and later dequantized to FP16 for inference.
 4.5.2. Learned ID Embeddings
-As discussed in Section 4.1, our pin embedder includes an ID embedding table.
-Effectively, it is represented as a large in-memory table that maps each pin ID’s hashes to an embedding, followed by some MLP layers.
-As this ID embedder takes up significant space in memory, especially when replicated, it is separated out from the main PinRec model and runs in its own CPU memory-optimized service.
-It is served with torchscript to limit latencies to single-digit milliseconds.
-4.4.3. Autoregressive Item Generation
-To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer with a KV Cache.
-Depending upon whether the model is outcome-conditioned or multi-token, we are able to sample or generate multiple representations respectively.
-After autoregressive generation, we use Faiss to retrieve a set of recommended items of size $N$ corresponding to the generated representations.
-Several optimizations are required for efficient inference:
+As discussed in Section 4.2, our pin embedder includes an ID embedding table that maps each pin ID’s hash to an embedding, followed by MLP layers.
+Since this ID embedder requires significant memory, especially when replicated, it runs as a separate CPU memory-optimized service to avoid competing with the main model for GPU resources.
+It is served with torchscript to achieve single-digit millisecond latencies.
+4.5.3. Autoregressive Inference
+We serve the main PinRec model for autoregressive generation on NVIDIA L40S GPUs.
+To generate an ordered set of item recommendations efficiently, we employ several optimizations for latency and cost efficiency:
 CUDA Graphs.
@@
 On subsequent decode steps, these are reused, and only new sequence elements are processed to produce $K$ and $V$ for the additional sequence dimension.
-Compression.
-Given that we generate multiple output representations with autoregressive generation, we observe there are often embeddings that are similar to one another.
-During Faiss retrieval, these are likely to collapse into a similar set of output items.
-To optimize this, we “compress” each generated embedding into a prior uncompressed embedding, in the order it is generated, if it is more similar than a given cosine similarity threshold.
-Compression significantly helps reduce the median number of embeddings generated by each model, with an average ratio of 6:1 for an unconditioned PinRec model and a smaller ratio of 3:1 for the multi-token, outcome-conditioned variant.
-Since we wish to fetch $N$ total items, we allocate these proportionally to each compressed embedding based on the total number of embeddings that were compressed into it.
-In the case of outcome-conditioning, we must perform this allocation independently for each action, in adherence with the chosen budget $\mathcal{B}$.
-The set of compressed embeddings alongside the allocations are passed to the Faiss retrieval stage.
 4.5.4. Faiss Retrieval
@@
 (b) Results for the user sequence specified in Subfigure 4(a).
-Figure 3. Item recommendations from PinRec-OC for a user history with cat-related search queries and pin interactions. History in reverse chronological order. Additional examples in Section A.2.
-Budget is: {repin: 0.4, outbound_click: 0.1, grid_click: 0.4, long_grid_click: 0.05, other: 0.05}. Generation runs for 6 steps (1 prefill, 5 decode).
+Figure 4.
+Item recommendations from PinRec-OC for a user history with cat-related search queries and pin interactions (history shown in reverse chronological order).
+Additional examples are provided in Section A.2.
 Our nearest-neighbor retrieval involves a Faiss IVF-HNSW index running on CPU hosts, with inner product as the similarity metric.
 As with the main model, this is also served with NVIDIA Triton with batching enabled.
-We run $k$NN retrieval to fetch the allocated number of Pins for each embedding, with an overfetch factor for later de-duplication.
-Sample results are visualized for a manually specified user history in Figure 3.
+We run kkNN retrieval to fetch the allocated number of Pins for each embedding, with an overfetch factor for later de-duplication.
+Sample results from this complete serving pipeline are visualized for a manually specified user history in Figure 4.
 5. Offline Evaluation
-In this section, we dissect the performance of the proposed sequential recommendation system, PinRec,  on generative retrieval across a range of offline evaluations.
+In this section, we dissect the performance of the proposed generative recommendation system, PinRec, on generative retrieval across a range of offline evaluations.
+For the results presented, PinRec uses a 12-layer transformer model (12 attention heads, 768 hidden dimensions) with an 80M-row ID embedding table (256 total dimensions, 8 dimensions per hash).
 Our experimental setting includes interactions on three major surfaces for the Pinterest platform, namely Homefeed, Related Pins, and Search.
@@
 5.1. Offline Evaluation Metrics
-We construct the evaluation set using interactions for a set of users disjoint from the training set and evaluate whether the proposed technique can produce a representation to retrieve future item embeddings among a set of 1M random items (“index”).
-In such tasks, recall @ $k$ is a standard performance metric, defined as the proportion of target items correctly retrieved by the model within the top-$k$ ranked items across an index.
-However, standard recall assumes the existence of a single target per sample.
-As mentioned in Section 4.3.2, such an approach fails to capture the assumption of a lack of strict ordering between user activity within a small range of interactions $\Delta$.
-Further, when the goal of the retrieval model is candidate generation for a downstream ranking model, the task naturally has more than one unique plausible target, making standard recall less relevant.
-We address these shortcomings by evaluating model performance on an unordered generalization of recall @ $k$.
-Given the autoregressively predicted item representations across all users $\mathbf{\hat{I}}$, target item embeddings across all users $\mathbf{I}$, and random item embeddings $\mathbf{N}$.
-Differently from standard recall, for each user $u\in\mathcal{U}$, unordered recall compares a set of predicted $\mathbf{\hat{I}}_{u}$ with a set of targets $\mathbf{I}_{u}$ of fixed size $\Delta$, calculating the proportion of targets $\mathbf{I}_{u}$ where there exists some $\mathbf{\hat{I}}_{u}$ that is within the top-$k$ ranked items.
-We formalize the notion of unordered recall in Equation 6.
-(6) $\begin{split}\mathrm{ur}_{k}(\mathbf{\hat{I}},\mathbf{I},\mathbf{N})\coloneq\frac{1}{|\mathbf{\mathbf{\hat{I}}}|}\times\sum_{i=1}^{|\mathbf{\hat{I}}|}\mathbf{1}\left\{\min_{j,z\in[\Delta]}\left|\{\mathbf{n}\in\mathbf{N}\mid\mathbf{\hat{I}}_{ij}^{\top}\mathbf{n}\geq\mathbf{\hat{I}}_{ij}^{\top}\mathbf{I}_{iz})\}\right|\leq k\right\}\end{split}$
-Note that since we perform autoregressive generation without teacher forcing, the transformer is not conditioned on any of the targets at any point.
-Beyond unordered retrieval performance, we examine the diversity of the recommendations.
-After retrieving the top-$k$ predicted items for each user, we compute the proportion of unique items retrieved across the evaluation set.
-This allows us to quantify the degree to which learned representations are biased towards popular or frequently occurring Pins in the dataset.
+We construct the evaluation set using interactions from users disjoint from the training set and evaluate whether the proposed technique can produce representations that retrieve future item embeddings from a set of $1M$ random items (the ”index”).
+For such retrieval tasks, recall@$k$ is a standard performance metric, measuring the proportion of target items correctly retrieved within the top-$k$ ranked items.
+We evaluate model performance using a sequence generalization of recall@$k$ that accounts for the multi-embedding nature of our autoregressive generation.
+Unlike standard recall, which compares individual predictions to individual targets, our unordered recall metric compares sets of predictions to sets of targets for each user.
+Given output item representations $\hat{\mathbf{I}}_{u}=\{\hat{\mathbf{i}}_{u,1},\hat{\mathbf{i}}_{u,2},\ldots,\hat{\mathbf{i}}_{u,m}\}$ and target item embeddings $\mathbf{I}_{u}=\{\mathbf{i}_{u,1},\mathbf{i}_{u,2},\ldots,\mathbf{i}_{u,n}\}$ for user $u$, along with random negative embeddings $\mathbf{N}$, we calculate the proportion of target items for which at least one predicted embedding retrieves it within the top-$k$ results:
+$\begin{split}\mathrm{ur}_{k}(\mathbf{\hat{I}},\mathbf{I},\mathbf{N})\coloneq\frac{1}{|\mathbf{\mathbf{\hat{I}}}|}\times\sum_{i=1}^{|\mathbf{\hat{I}}|}\mathbf{1}\left\{\min_{j,z\in[\Delta]}\left|\{\mathbf{n}\in\mathbf{N}\mid\mathbf{\hat{I}}_{ij}^{\top}\mathbf{n}\geq\mathbf{\hat{I}}_{ij}^{\top}\mathbf{I}_{iz})\}\right|\leq k\right\}\end{split}$
+where $\text{rank}(\mathbf{i},\hat{\mathbf{i}},\mathbf{N})$ denotes the rank of target item $\mathbf{i}$ when items in $\mathbf{N}\cup{\mathbf{i}}$ are sorted by cosine similarity to prediction $\hat{\mathbf{i}}$.
+This metric captures whether any of the generated embeddings successfully retrieves each target item, reflecting the practical scenario where multiple predictions can contribute to successful recommendation.
+We perform autoregressive generation without teacher forcing, ensuring the transformer is not conditioned on any target items during prediction, making the evaluation more realistic for deployment scenarios.
+Beyond retrieval performance, we examine recommendation diversity by computing the proportion of unique items retrieved across all users in the evaluation set after selecting top-$k$ predictions.
+This quantifies whether learned representations exhibit bias toward popular items versus providing diverse recommendations across the user base.
 5.2. Offline Results
-In this section, we compare PinRec with an existing sequence-based user modeling representations, Pinnerformer (Pancha et al., 2022), SASRec (Kang and McAuley, 2018b), and TIGER (Rajput et al., 2023), in terms of offline evaluation metrics.
-Further, we quantify the impacts of our design on the performance and diversity of recommendations.
+In this section, we compare PinRec against existing sequence-based user modeling approaches, including Pinnerformer (Pinnerformer22), SASRec (kang2018self), and the leading generative recommendation method TIGER (Tiger23).
+We also evaluate our transformer architecture against the proposed HTSU (HSTU24) module.
+Furthermore, we quantify the impacts of our design choices on both recommendation performance and diversity.
 5.2.1. Comparison with Prior Method
-To validate that PinRec outperforms existing sequential techniques in terms of item recommendation in offline evaluation, we compare its performance to PinnerFormer.
 We evaluate three different variants of PinRec.
@@
 As in training (where we cannot modify the evaluation dataset post-hoc), we condition on the target item’s action for outcome-conditioned variants.
-Our comparison in Table 1 shows that PinRec outperforms a similarly sized Pinnerformer across the board.
+For simplicity, we reproduce results for the outcome-conditioned variant only when comparing against HTSU (additional details in Section A.1.1).
+The impact of negative sampling and model scaling effects are analyzed in Section A.1.2 and Section A.1.3, respectively.
+Our comparison in Table 1 shows that PinRec outperforms other baselines across the board.
 Importantly, outcome-conditioning demonstrates notable improvements over PinRec-UC (unconditioned) in unordered recall @ 10 (e.g., +2-3% on Homefeed and Related Pins).
@@
 Model Homefeed Related Pins Search
-TIGER (Rajput et al., 2023) 0.208 0.230 0.090
-SASRec (Kang and McAuley, 2018b) 0.382 0.426 0.142
-PinnerFormer (Pancha et al., 2022) 0.461 0.412 0.257
+SASRec (kang2018self) 0.382 0.426 0.142
+PinnerFormer (Pinnerformer22) 0.461 0.412 0.257
+TIGER (Tiger23) 0.208 0.230 0.090
+HTSU-OC (HSTU24) 0.596 0.539 0.179
 PinRec-UC 0.608 0.521 0.350
@@
 5.2.2. Outcome-Conditioned Generation
-We seek to understand whether we can achieve “controllability” between the desired action and true action; for instance, does conditioning on a “repin” generate an output representation that truly aligns with content that will be repinned?
-If we are able to control the outcome through our item recommendations, then we will observe that for the specific action that we condition on, there are significant gains in offline recall.
-Conversely, we expect that the recall for items corresponding to other actions will decrease since our action-conditioned generation should ideally induce only that chosen action.
-As a baseline, we compare to the unconditioned variant of PinRec, which does not take into account the desired action.
-For simplicity, we allocate the entire budget entirely to one action and perform offline evaluation conditioning across each action.
+We assess whether conditioning on a desired action (e.g., “repin”) leads to outputs that elicit that action.
+If our recommendations successfully control outcomes, we expect higher offline recall for the conditioned action and lower recall for others.
+As a baseline, we compare against the unconditioned PinRec variant.
+For simplicity, we allocate the entire budget to one action and perform offline evaluation conditioning across each action.
 We show the result of this analysis in Figure 5.
@@
 5.2.3. Utility of Multi-Token Generation
-Beyond the improvement in performance, we evaluate the benefits of multi-token generation in terms of efficiency, diversity, and performance of generations relative to the number of chosen generations per timestep.
-We examine the tradeoff between generating more representations per timestep versus longer autoregressive generation, while keeping the number of output embeddings constant.
+We evaluate the benefits of multi-token generation as a function of the number of generations produced per timestep.
+Specifically, we analyze the tradeoff between generating more representations in fewer steps versus fewer representations over more steps, while keeping the total number of output embeddings constant.
 Since the strongest offline performance aside from the multi-token variant is by PinRec-OC, we consider it as a baseline.
@@
 While next-token generation (i.e., 1 embedding per step) is the most inefficient and yields the poorest performance (where we perform similarly to PinRec-OC), we observe significant benefits across the board in increasing the number of embeddings per step.
-At 16 generations per step, we demonstrate $\sim$10x reduction in latency over PinRec-OC, with improved unordered recall (+16.0%) and diversity (+21.3%).
+At 16 generations per step, we demonstrate a $\sim$10x reduction in latency over PinRec-OC, with improved unordered recall (+16.0%) and diversity (+21.3%).
 Figure 6. Tradeoff between unordered recall @ 10 (top), proportion unique (bottom), and generation latency (right) with the number of multi-token generations per step.
@@
 The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other candidate generators, to produce the final results shown to users.
-Since we employ a heterogeneous sequence of Pins and search queries, we focus on the Homefeed and Search surfaces.
-On these surfaces, we provide personalized item recommendations through PinRec, leading to significant lifts in action metrics, session metrics, and intent fulfillment.
+We tested on the three major surfaces: Homefeed, Search and Related Pins.
 6.1. Homefeed
@@
 6.1.2. Metrics
-Pinterest uses “session fulfillment” as an informative and holistic whole-funnel metric to capture (a) the upper funnel of user intent expression (e.g., when a user clicks on a pin in the grid), (b) the middle funnel of curating Pins (e.g., repin), and (c) the lower funnel of realizing intent (e.g., clicking on a product pin to visit its website).
-Any of the aforementioned actions results in a session being marked as “fulfilled”, whereas the absence of these results in an “unfulfilled” session.
+We use engagement actions such as repins, grid clicks, and time spent.
+Additionally, we count “fulfilled sessions,” where the user takes at least one action in the session, and “unfulfilled sessions,” where the user does not take any action.
 6.1.3. Results
 In Figure 7, we visualize a Pareto front of the repin rate (relative to impressions) and grid click rate for the PinRec candidate generator by itself, with absolute numbers hidden due to non-public information.
-By reducing the grid click budget and correspondingly increasing the repin budget, we see the expected result as the grid click rate drops and repin rate increases (i.e., we demonstrate online “controllability”).
-Similarly, we observe a significant correlation between the budgets and the overall A/B experiment metric lifts for Homefeed grid clicks and repins in Figure 7.
-Further, we note that across both variants in Figure 7, for $\mathcal{B}(\text{repin})\geq 0.6$, there are statistically significant overall online lifts in Homefeed repin.
-We achieve statistically significant lifts in Homefeed grid click for all budget groups on each variant.
+By reducing the grid click budget and correspondingly increasing the repin budget, we see the expected result as the grid click rate drops and repin rate increases.
+Similarly, we observe a significant correlation between the budgets and the overall A/B experiment metric lifts for Homefeed grid clicks and repins in Figure 8 in Appendix.
 (a) PinRec-OC
@@
 Figure 7. Visualization of Pareto fronts for online rate of grid click versus repin for items recommended only by PinRec variants, with error bars denoted in grey.
-(a) PinRec-OC
-(b) PinRec-{MT, OC}
-Figure 7. Visualization of Pareto fronts for online lift % in Homefeed grid clicks versus repin for PinRec variants.
-To examine the holistic impact of our candidate generators, we examine the lift in fulfilled sessions for each PinRec variant, selecting the group which maximizes this metric.
-To demonstrate notable site-wide, session-level, and global metrics improvements, we present the CUPED-adjusted  (Deng et al., 2013) percentage lift in metrics across Homefeed and the entire site (Table 2) for the groups that maximize fulfilled sessions.
-Relative to the control group, the unconditioned variant of PinRec (PinRec-UC) demonstrates improvements in unfulfilled sessions and in Homefeed and site-wide grid clicks.
-In addition to those lifts, the outcome-conditioned variant (PinRec-OC) shows significant lift in site-wide fulfilled sessions.
-The multi-token variant, PinRec-{MT, OC}, significantly increased successful site sessions, time spent on site, and site-wide grid clicks, which we believe represent the most comprehensive site-wide improvements.
-Table 2. CUPED-adjusted metrics improvements from online A/B experiments for PinRec variants as candidate generators in Homefeed.
+For each PinRec variant, we selected the variant with the best fulfilled sessions and measured the percentage lift in metrics across Homefeed and the entire site (Table 2).
+Relative to the control group, the unconditioned variant of PinRec, PinRec-UC, demonstrates improvements in unfulfilled sessions and in Homefeed and site-wide grid clicks.
+In addition to those lifts, the outcome-conditioned variant, PinRec-OC, shows significant lift in site-wide fulfilled sessions.
+The multi-token variant, PinRec-{MT, OC}, significantly increased successful site sessions, time spent on site, and site-wide grid clicks.
+Table 2. A/B experiments for PinRec variants as candidate generators in Homefeed.
 Type Metric UC OC {MT, OC}
@@
 On the Search surface, we evaluate the unconditioned variant of PinRec.
-As shown in Table 3, PinRec-UC improves search metrics significantly relative to a control group.
-We observe session-wide search “fulfillment” increase, i.e., when a search session results in positive interaction such as repins or long outbound clicks.
-Table 3. CUPED-adjusted metrics improvements from online A/B experiments for PinRec as candidate generator in Search retrieval.
+As shown in Table 3, PinRec-UC improves search metrics significantly relative to a control group, with increased session-wide search fulfillment (sessions resulting in positive interactions like repins or long clicks).
+When deployed for promoted product retrieval, the same embeddings achieve a 1.83% reduction in CPA and 1.87% increase in shopping conversion volume in the US, indicating improved relevance and user engagement with sponsored content.
+Table 3. A/B experiments for PinRec as candidate generator in Search retrieval.
 Metric Type Metric PinRec-UC
-Overall Metrics Search Fulfillment Rate +0.48%
-Action Metrics Search Repins +0.84%
-Search Grid Clicks +1.46%
+Overall Metrics Search Fulfillment Rate +2.27%
+Action Metrics Search Repins +4.21%
+Search Grid Clicks +3.00%
+6.3. Related Pins
+On the Related Pins surface, we evaluate the outcome-conditioned variant of PinRec.
+As shown in Table 4, PinRec-OC improves Related Pins metrics significantly relative to a control group.
+We observe session-wide increase in fulfilled sessions and increase in sitewide engagement.
+Table 4. A/B experiments for PinRec variants as candidate generators in Related Pins.
+Type Metric PinRec-OC
+Overall Fulfilled Sessions +0.26%
+Metrics Time Spent +0.30%
+Unfulfilled Sessions -1.07%
+Action Site-wide Grid Clicks +0.50%
+Metrics Sitewide Repins +0.63%
 7. Conclusion
-We introduced PinRec, a generative retrieval system developed for Pinterest’s Homefeed and Search feed.
+We introduced PinRec, a generative retrieval system developed for Pinterest’s Homefeed, Search and Related Pins feed.
 To the best of our knowledge, this paper represents the first comprehensive study on deploying generative retrieval in a web-scale application.
-The incorporation of two key innovations — outcome-conditioned generation and multi-token inference — enabled PinRec to enhance several business metrics without substantial cost increases.
+The incorporation of two key innovations — outcome-conditioned generation and temporal multi-token inference — enabled PinRec to enhance several business metrics without substantial cost increases.
 Several promising future directions can be pursued.
 First, applying generative approaches to ranking, as explored by Zhai et al.
-(Zhai et al., 2024), could be beneficial.
+(HSTU24), could be beneficial.
 Second, optimizing input sequences to the model, akin to prompt engineering in large language models, is another avenue worth exploring.
 Third, developing language model alignment techniques similar to Ouyang et al.
-(Ouyang et al., 2022) could enhance the use of engagement signals from users.
+(PPO22) could enhance the use of engagement signals from users.
 Overall, this paper highlights a promising avenue for utilizing generative retrieval in large-scale recommender systems.
-Acknowledgements.
-We thank collaborators such as Deepak Agarwal, Divyansh Agarwal, Edoardo Botta, Bee-Chung Chen, Bowen Deng, Pong Eksombatchai, Kurchi Subhra Hazra, Se Won Jang, Saurabh Vishwas Joshi, Krishna Kamath, Sujay Khandagale, James Li, Cristian Lopez, Matt Madrigal, Lei Pan, and Ruijia Wang for their critical contributions to our work and for facilitating productive discussions.
 Appendix A Appendix
@@
 Specifically, outcome conditioning and multi-token prediction can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements.
-To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (Zhai et al., 2024), as opposed to the multi-head attention-based causal transformer.
-Note that we do not leverage the ranking task or any sequence sampling explored in (Zhai et al., 2024) as it is not relevant for our specific use case.
+To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (HSTU24), as opposed to the multi-head attention-based causal transformer.
+Note that we do not leverage the ranking task or any sequence sampling explored in (HSTU24) as it is not relevant for our specific use case.
 For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
@@
 A.1.2. Ablating Negative Sampling
-Following from Pancha et al.
-(2022), we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function.
-Though we ultimately end up with similar conclusions as Pancha et al.
-(2022), we show these findings for maximum clarity regarding our design choices.
+Following from Pinnerformer22, we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function.
+Though we ultimately end up with similar conclusions as Pinnerformer22, we show these findings for maximum clarity regarding our design choices.
 Random Negatives
@@
 Hence, we opt to not use larger sized transformers.
-Vocabulary Parameters
+Vocabulary Parameters.
 As presented in the main text, the ID embeddings consist of roughly 10 billion parameters.
@@
 Table 10. Online experimental results for different user cohorts and tail item metrics, across PinRec-UC and PinRec-OC.
-In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 9(a).
+In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 10(a).
 Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users.
@@
 We believe the evidence suggests that our generative retrieval solution is equal to or better across inactive/new users and tail items relative to existing control, and perhaps more importantly, it shows that outcome conditioning can help in boosting performance in these criteria, which is one of its goals.
+(a) PinRec-OC
+(b) PinRec-{MT, OC}
+Figure 8. Visualization of Pareto fronts for online lift % in Homefeed grid clicks versus repins for PinRec variants.
 A.1.5. Serving Efficiency
@@
 A.2. Visualization of Multi-Token Impact
-In this section, we present some additional qualitative examples through visualizations of the retrieved examples for some sample user histories across the windowed multi-token and non-multi-token variants.
+(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
+(b) Results for PinRec-OC candidate generator for user history in subfigure above. Notably only includes Pins related to nails, the user’s most recent interaction.
+(c) Results for PinRec-{MT, OC} candidate generator for user history in subfigure above. Includes a mix of Pins related to both tattoos and nails.
+Figure 9. Comparison of results for a manually-specified user history.
+(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
+(b) Results for PinRec-OC candidate generator for given user history. Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
+(c) Results for PinRec-{MT, OC} candidate generator for given user history. Includes a mix of Pins related to nurseries and food.
+Figure 10. Comparison of results for another manually-specified user history.
+In this section, we present additional qualitative examples through visualizations of retrieved examples for sample user histories, comparing multi-token and non-multi-token variants.
 Through these visualizations, we are able to demonstrate a reasonable and consistent qualitative difference in the types of pins retrieved, justifying the improved session-wide and sitewide online metrics for PinRec-{MT, OC}.
@@
 These are critical qualities for a platform like Pinterest.
-(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
-(b) Results for PinRec-OC candidate generator for user history in subfigure above. Notably only includes Pins related to nails, the user’s most recent interaction.
-(c) Results for PinRec-{MT, OC} candidate generator for user history in subfigure above. Includes a mix of Pins related to both tattoos and nails.
-Figure 8. Comparison of results for a manually-specified user history.
-(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
-(b) Results for PinRec-OC candidate generator for given user history. Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
-(c) Results for PinRec-{MT, OC} candidate generator for given user history. Includes a mix of Pins related to nurseries and food.
-Figure 9. Comparison of results for another manually-specified user history.

### v5

@@
-PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems
+PinRec: Unified Generative Retrieval for Pinterest Recommender Systems
 Abstract.
-Generative retrieval methods utilize generative sequential modeling techniques, such as transformers, to generate candidate items for recommender systems.
-These methods have demonstrated promising results in academic benchmarks, surpassing traditional retrieval models like two-tower architectures.
-However, current generative retrieval methods lack the scalability required for industrial recommender systems, and they are insufficiently flexible to satisfy the multiple metric requirements of modern systems.
-This paper introduces PinRec, a novel generative retrieval model developed for applications at Pinterest.
-PinRec utilizes outcome-conditioned generation, enabling modelers to specify how to balance various outcome metrics, such as the number of saves and clicks, to effectively align with business goals and user exploration.
-Additionally, PinRec incorporates multi-token generation to enhance output diversity while optimizing generation.
-Our experiments demonstrate that PinRec can successfully balance performance, diversity, and efficiency, delivering a significant positive impact such as +2% sitewide clicks and +4% search repins.
-This paper marks a significant milestone in generative retrieval, as it presents, to our knowledge, the first rigorous study on implementing generative retrieval at the scale of Pinterest.
-generative retrieval, recommender systems, transformer, item recommendation
+Generative retrieval methods employ sequential modeling techniques, like transformers, to generate candidate items for recommender systems.
+These methods have demonstrated promising results in academic benchmarks, surpassing traditional retrieval models such as two-tower architectures.
+However, a key limitation is that current approaches require a separate model for each product surface, as building a unified model that accommodates the different business needs of various surfaces has proven challenging.
+Furthermore, existing methods often fail to capture the evolution of user interests over a sequence, focusing instead on only predicting the next item.
+This paper introduces PinRec, a novel unified generative retrieval model for all of Pinterest’s recommendation surfaces, including home feed, search, and related pins.
+PinRec is pretrained on user activity sequences aggregated across surfaces, then fine-tuned for each surface using that surface’s impression data.
+This pretraining–fine-tuning approach enables a single unified model while still adapting to the needs of individual surfaces.
+To better align recommendations with surface-specific business goals, PinRec incorporates a novel outcome-conditioned generation mechanism that targets different outcomes for each surface, which further enhances the impact of fine-tuning.
+Our experiments show that PinRec balances performance, diversity, and efficiency, delivering significant gains such as +4% increase in search saves.
+To our knowledge, this paper presents the first rigorous study of a unified generative retrieval model built and deployed at Pinterest scale, marking a significant milestone in the field.
+generative retrieval, recommender systems
 journalyear: 2025copyright: ccccs: Information systems Retrieval models and rankingccs: Information systems Personalization
 Figure 1.
-Illustration of PinRec, a generative item retrieval technique for heterogeneous user journeys on Pinterest.
-Sequences of user searches, engagements, and outcome-conditioning (bottom) are used to recommend Pins (top).
+Illustration of PinRec, a unified generative retrieval model for heterogeneous user journeys on Pinterest.
+Sequences of user searches and activities are used to recommend Pins (top).
+PinRec powers retrieval in multiple surfaces optimizing surfaces’ unique business metrics like Saves for Home Feed and Product Click for Search
 1. Introduction
-Pinterest is a visual discovery platform where over 550 million monthly active users search for, save, and shop for ideas, represented by Pins (pins).
-Given the massive volume of Pins — for instance, users save over 1.5 billion Pins per week (pins2) — it is essential to provide personalized recommendations in Pinterest feeds and search results.
+Pinterest is a visual discovery platform where over 600 million monthly active users search for, save, and shop for ideas, represented by Pins (Pinterest, 2025).
+Given the massive volume of Pins—for instance, users save over 1.5 billion Pins per week (Pinterest, 2024)—it is essential to provide personalized recommendations in Pinterest feeds and search results.
 To achieve this, recommender systems often leverage a two-stage approach: retrieval and ranking.
-Retrieval models select thousands of potentially relevant candidate items, then ranking models prioritize the top items to display to the user.
-Since ranking models depend on the retrieval results, it is important for retrieval models to identify diverse, fresh, and relevant items.
+Retrieval models select thousands of potentially relevant candidate Pins, then ranking models prioritize the top Pins to display to the user.
+Since ranking models depend on the retrieval results, it is important for retrieval models to identify diverse, fresh, and relevant Pins.
 Traditionally, retrieval methods employ “two-tower” models.
-The query tower computes query embeddings, while the item tower computes item embeddings.
-These models retrieve items with the highest similarity scores given a query embedding.
-As an alternative to two-tower models, recent work has proposed generative retrieval.
-Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (vaswani2017attention), and then use those tokens to retrieve items.
-Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (KangM18; Recformer23; UnifyingGenRetrieval24).
-Using generative retrieval models for large scale recommender systems poses many challenges at the scale of Pinterest.
-Firstly, the generative model should be capable of efficiently optimizing for different engagement outcomes such as maximizing social actions like saves and comments versus passive engagements like clicks, because business needs are constantly evolving to optimize combinations of different metrics.
-Secondly, the generated item sequences must possess sufficient diversity to facilitate users’ exploration of varied ideas; one-step generation of a single item representation may not suffice.
-Thirdly, given their increased complexity and sequential generation, generating a set of recommended items requires higher computational cost than computing an embedding with two-tower models.
-To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges.
-Most methods developed within academic settings (KangM18; UniS4Rec22; Recformer23) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic.
-While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (HSTU24; Tiger23) and are not capable of optimizing multiple target outcomes.
-Moreover, these methods do not address the challenge of generating a sequence of diverse items.
-Because generating multiple items sequentially is computationally expensive, existing approaches typically generate only the next item rather than a full set at once.
-Figure 2. Modeling diagram of PinRec
-In this paper, we present PinRec, a generative retrieval technique for real-world recommender systems (Figure 2).
-PinRec builds on other generative retrieval models, leveraging transformers to perform next token prediction (KangM18; HSTU24; Tiger23).
-Unlike previous methods, however, PinRec addresses and resolves three aforementioned challenges for generation of optimal recommendations: multi-objective generative capabilities, computational efficiency, and item diversity.
-Our main contributions in the paper are summarized below.
-  • Outcome-Conditioned Generative Retrieval: PinRec employs outcome-conditioned generation, enabling the optimization of multiple business metrics.
-For each outcome (e.g., business metrics such as user saves or clicks), PinRec generates the next item token that is most likely to maximize that specific outcome, and then retrieves candidate items to balance these objectives according to the desired trade-offs (Fig.
+The query tower computes query embeddings, while the item tower computes Pin embeddings.
+These models retrieve Pins with the highest similarity scores given a query embedding.
+As an alternative to two-tower models, recent work has proposed Generative Retrieval (GR).
+GR models leverage a sequence of tokens using generative models, such as Transformers (Vaswani et al., 2017), and then use those tokens to retrieve Pins.
+GR has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018a; Li et al., 2023; Yang et al., 2024).
+In this paper, we aim to develop a unified GR model that can serve multiple Pinterest recommendation surfaces (e.g., search, home feed, related Pins) by being fine-tuned on the target surface.
+Because GR models operate over a single token sequence, they can combine heterogeneous inputs—such as query text and user history—into one representation, enabling parameter sharing and cross-surface transfer.
+Moreover, the training signal is naturally cross-surface: interaction logs from different surfaces can be expressed in a common next-item (or next-token) prediction framework, allowing the model to learn from a unified mixture of surface-specific trajectories.
+In contrast, two-tower systems often require surface-specific query towers, objectives, and/or retrieval indices, which increases complexity and can limit generalization across surfaces.
+Despite these advantages, translating a unified GR approach into a production retriever that serves multiple Pinterest surfaces introduces several practical challenges.
+First, it must efficiently optimize surface-specific objectives, since each surface defines success differently: search emphasizes outbound clicks that reflect offsite conversion intent, related Pins emphasizes session depth and continued exploration, and home feed emphasizes saves and discovery (clicks and downstream exploration).
+This is further complicated when key metrics are sparse—for example, outbound product clicks in search are relatively rare events.
+Second, user behavior and demographics differ by surface: upstream surfaces at the start of a session (e.g., the home feed) have a higher share of less-active users, whereas downstream surfaces (e.g., related Pins) skew toward more active users who take stronger actions.
+Third, user intent can evolve even within a single session, and a generative retriever must adapt its generated sequence accordingly.
+Fourth, due to sequential decoding, GR models typically incur higher computational cost and latency than two-tower models, which only require embedding computation and nearest-neighbor search.
+Figure 2.
+PinRec is pretrained with the user activity sequences consisting of items (Pins) and search queries.
+PinRec is then fine-tuned for each surface specific data that contains user actions and impressed items (Pins).
+To our knowledge, no prior work addresses all of these requirements in a single unified GR model, namely (i) surface-specific success metrics, (ii) heterogeneity in cross-surface data distributions, (iii) in-session user intent shifts, and (iv) the latency/cost constraints.
+Recent industrial GR systems such as TIGER and HSTU (Zhai et al., 2024; Rajput et al., 2023) are designed for a single next-token/next-item prediction objective, and in practice are typically trained against a unified engagement target (e.g., any action) rather than explicitly aligning the learning signal to surface-specific goals (e.g., outbound clicks for search, saves for home feed).
+As a result, fine-tuning can yield limited gains when the surface “true north” metrics differ (see Sec.
+5.2.3).
+Separately, foundational sequence models like RecFormer and PLUM can be fine-tuned for individual applications (Hou et al., 2022; Li et al., 2023; He et al., 2025) and can better match surface-specific user/item distributions, but they are often prohibitively expensive at Pinterest scale due to their reliance on LLM backbones; moreover, they are trained for a single generic engagement target, leaving the objective-misalignment issue unchanged.
+Finally, existing GR approaches do not support multi-step generation at retrieval time; because autoregressive decoding is costly, they commonly generate only the next-item candidates, limiting their ability to produce a longer candidate sequence that anticipates or tracks within-session shifts in intent.
+Figure 3. Generation phase of PinRec
+In this paper, we present PinRec, a foundational GR model for Pinterest recommendation surfaces.
+We pretrain PinRec on user activity data aggregated across surfaces (Chen et al., 2025), and then fine-tune it on surface-specific data to align the model with each surface’s business goals (Fig.
 2).
-Through both offline and online evaluations, we show that tailoring item recommendations to particular outcomes enables designers to effectively achieve those specified objectives.
-  • Multi-Token Generation for Efficient Sequence Generation: One obstacle to generating an entire sequence of items for retrieval is that traditional models typically produce only one item at a time.
-To address this, we propose a multi-token generation framework (Multitoken24), in which PinRec generates multiple future items in parallel, as illustrated in Figure 2.
-This approach significantly reduces the number of steps needed to produce diverse output embeddings, resulting in substantial offline and online improvements in performance, diversity, and efficiency.
-  • Efficient Serving: We detail how we optimize autoregressive generation for optimal item recommendation, implement the serving infrastructure, and allow for horizontal scaling to serve hundreds of millions of users.
-To our knowledge, this study is the first to disclose the details of infrastructure optimization for industry-scale generative retrieval systems.
+Unlike prior GR methods, PinRec addresses four aforementioned challenges: (1) optimizing surface-specific objectives, (2) adapting to each surface’s data distribution, (3) capturing evolving user intent within a session via sequence generation, and (4) doing so with only a marginal increase in serving cost.
+We summarize our solutions to each challenge below.
+  • Outcome-Conditioned Generative Retrieval: To optimize surface-specific objectives, PinRec employs outcome-conditioned generation, enabling the optimization of multiple business metrics.
+For each outcome (e.g., user saves or clicks), PinRec generates a sequence of Pin tokens conditioned on that outcome and retrieves candidate Pins; the generated sequences can then be combined to balance objectives according to desired trade-offs (Fig.
+3).
+Our offline and online evaluations show that outcome conditioning enables PinRec to adapt to the distinct business needs of each surface.
+  • Fine-tuning for each surface: To adapt for each surface, we employ a pretraining–fine-tuning framework.
+We first pretrain the base PinRec model on user activity sequences aggregated across all Pinterest surfaces.
+To tailor the model to each surface’s user population and behavior patterns, we then fine-tune PinRec on surface-specific impression logs.
+Using this approach, we deployed PinRec in production across three surfaces.
+To our knowledge, this is the first paper to apply a pretraining–fine-tuning paradigm for generative retrieval across multiple recommendation surfaces at web scale.
+  • Efficient Serving with Auto-regressive generation: To capture evolving user interest, PinRec generates item sequences auto-regressively.
+To mitigate the cost of generation, we describe optimizations that enable low-latency decoding, detail the serving infrastructure, and show how we scale the system horizontally to serve hundreds of millions of users.
+To our knowledge, this is the first study to disclose infrastructure optimizations for an industry-scale generative retrieval system.
 We will make the model code publicly available upon publication.
 2. Related Work
-Sequential Recommendation Systems: Sequential recommendation systems operate on a sequences of items with which a user has interacted (KangM18).
-Earlier work in academia has used recurrent neural networks (RNN) (GRU4REC16), but later work has often adopted transformers (KangM18; BERT4REC19).
-Recent studies have harnessed the extensive knowledge of pre-trained language models (UniS4Rec22; Recformer23).
-In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pinnerformer22).
-The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models) (Spotify20; rethink; Transact23; TWINV224).
-Another approach in sequential recommendation systems is the generative approach, where the model learns to generate the items to recommend.
-Tiger23 used transformers to generate items by producing a sequence of semantic ID tokens (RQVAE22; SemanticIDGeneral24).
-HSTU24 also proposed a generative retrieval method.
-However, these studies have mainly focused on generating a single item to optimize a single action.
-In contrast, our research aims to generate sequences of multiple items to optimize for multiple outcomes.
+Sequential Recommendation Systems: Sequential recommendation systems operate on a sequences of items with which a user has interacted (Kang and McAuley, 2018a).
+Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018a; Sun et al., 2019).
+Recent studies have harnessed the extensive knowledge of pre-trained language models (Hou et al., 2022; Li et al., 2023; He et al., 2025).
+In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pancha et al., 2022).
+The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models) (Hansen et al., 2020; Xu et al., 2022; Xia et al., 2023; Si et al., 2024).
+A different line of work in sequential recommendation is generative recommendation, where the model directly generates the items to recommend.
+Rajput et al.
+(2023) uses a Transformer to generate items by producing sequences of semantic ID tokens (Lee et al., 2022; Singh et al., 2024), while Zhai et al.
+(2024) proposes a new architecture (HSTU).
+However, these early methods have largely been evaluated in single-application settings; extending them to multiple surfaces is challenging because they typically optimize only a single objective (e.g., whether an action occurs).
+Very recently, Deng et al.
+(2025); Dai et al.
+(2025) proposed multi-stage training frameworks that unify retrieval and ranking within a single surface.
+This represents vertical unification, integrating different stages of the recommendation funnel for one surface.
+In contrast, we pursue horizontal unification: we fine-tune a single retrieval model to serve multiple surfaces.
+Both directions are important, but we expect horizontal unification to be easier to translate into product impact, since the unified model can be integrated into existing multi-stage funnels as a drop-in replacement at the retrieval stage.
 Multi-Objective Generation: We discuss generative models that consider multiple objectives.
-While RL algorithms have shown this capability in industry (SelfsupervisedRL20), studies in language modeling suggest RL often needs extensive tuning for optimal generation (PPO22).
-Recent research has introduced more streamlined and scalable techniques based on supervised learning, such as goal-conditioned methods (IterativeSupervised21) and prompting strategies (PromptingRL22).
+While RL algorithms have shown this capability in industry (Xin et al., 2020), studies in language modeling suggest RL often needs extensive tuning for optimal generation (Ouyang et al., 2022).
+Recent research has introduced more streamlined and scalable techniques based on supervised learning, such as goal-conditioned methods (Ghosh et al., 2021) and prompting strategies (Xin et al., 2022).
 PinRec builds on these approaches to enhance scalability.
@@
 For each user $u$, we maintain their interaction history $H(u,t_{\text{max}})=\langle(i_{1},a_{1},s_{1},t_{1}),(i_{2},a_{2},s_{2},t_{2}),\ldots,(i_{m},a_{m},s_{m},t_{m})\rangle$, where each tuple $(i_{j},a_{j},s_{j},t_{j})$ represents an interaction with item $i_{j}\in\mathcal{I}$ via action $a_{j}\in\mathcal{A}$ on surface $s_{j}\in\mathcal{S}$ at timestep $t_{j}\leq t_{\text{max}}$.
-The sequence is chronologically ordered such that $t_{1}\leq t_{2}\leq\ldots\leq t_{m}$.
+The sequence is chronologically ordered, $t_{1}\leq t_{2}\leq\ldots\leq t_{m}$.
+Surface-specific Feed View.
+When user $u$ visits surface $s$ at time $t$, we observe a feed-view event $F(u,s,t)=\langle(i_{1},a_{1},t_{1}),\ldots,(i_{m},a_{m},t_{m})\rangle,$ where each tuple $(i_{j},a_{j},t_{j})$ indicates that item $i_{j}\in\mathcal{I}$ was shown to the user at $j$-th position, and the user’s action on that impression is encoded by $a_{j}\in\mathcal{A}\cup\{\text{imp}\},$ with imp denoting an impression with no explicit action taken (i.e., the item was shown but not clicked/saved/etc.).
+If an explicit engagement occurs (e.g., click, save), then $a_{j}\in\mathcal{A}$.
+The timestamp $t_{j}$ is the time of action when $a_{j}\in\mathcal{A}$; otherwise (no action), we set $t_{j}=t$.
 Sequential Recommendation Model.
@@
 In this section, we describe the model architecture and serving infrastructure for PinRec.
-We start with an overview of the model, including outcome-conditioned generation and multi-token prediction.
-Next, we explain how items in the sequence are encoded and processed, and how the training loss is computed.
+We start with an overview of the model, including outcome-conditioned generation.
+Next, we explain how items in the sequence are encoded and processed.
+Then we talk about how we perform pretraining on a unified data and how we fine-tune for a surface-specific data.
 Finally, we describe inference mode operation and provide details about our real-time serving architecture, which is powered by Nvidia Triton.
@@
 Outcome-Conditioned Generation.
-Unlike language modeling, where the output sequence receives a single reward value (PPO22), users interact with recommended items through diverse engagement types—saving, sharing, commenting, or clicking.
-To effectively optimize across these varied engagement metrics, generative retrieval methods must account for this behavioral diversity (MultiBehaviorSASRec24).
+Unlike language modeling, where the output sequence receives a single reward value (Ouyang et al., 2022), users interact with recommended items through diverse engagement types—saving, sharing, commenting, or clicking.
+To effectively optimize across these varied engagement metrics, generative retrieval methods must account for this behavioral diversity (Liu et al., 2024).
 Furthermore, practical deployment often requires specifying the target surface or context for the next recommendation, even when the user’s historical sequence spans multiple surfaces.
@@
 where $\mathbf{h}_{u,t}$ is the hidden state at timestep $t$, $\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k}$ are the conditioning embeddings for different outcome types (e.g., $\mathbf{c}_{1}=\mathbf{e}_{\text{action}}$, $\mathbf{c}_{2}=\mathbf{e}_{\text{surface}}$), and $\hat{\mathbf{i}}_{u,t}$ is the resulting conditioned item representation.
-Temporal Multi-token Prediction.
-Multi-token prediction in language models involves predicting multiple successive tokens at positions $t+1$ to $t+k$ given the current position $t$ (gloeckle2024better).
-This is typically implemented using separate output heads for each future position.
-However, in recommender systems, previous research (Pinnerformer22) suggests that the difference in time between actions is much more important than the position index alone, because a single position increment can correspond to very different time intervals for active users versus dormant users.
-To address this, we propose a time-based multi-token approach, in which tokens are generated for each time interval $\delta$ (e.g., generating recommendations for the user for the present, 30 seconds later, 1 minute later, and so on).
-Specifically, we condition our output head on the relative temporal distance from the current action, enabling a single parameterized head to generate predictions at various future time intervals:
-(2) $\hat{\mathbf{i}}_{u,t+\delta}=O(\mathbf{h}_{u,t},\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{k},\mathbf{e}_{\delta})$
-where $\mathbf{e}_{\delta}$ is the learnable embedding for temporal offset $\delta\in\{1,2,\ldots,k\}$, and $\hat{\mathbf{i}}_{u,t+\delta}$ represents the predicted item at timestep $t+\delta$.
 4.2. Input Sequence Representation
 Given the heterogeneous nature of user interactions $(i_{j},a_{j},s_{j},t_{j})$ in our interaction history $H(u,t_{\text{max}})$, we need to transform each element into appropriate representations for the transformer architecture.
-A crucial design decision is defining representations for heterogeneous item types.
-There are two main ways to represent items in a sequence: (1) by using dense embedding vectors (HSTU24; KangM18), and (2) by learning a sequence of discrete IDs for each item, known as Semantic IDs (Tiger23; SemanticIDGeneral24).
+There are two main ways to represent items in a sequence: (1) by using dense embedding vectors (Zhai et al., 2024; Kang and McAuley, 2018a), and (2) by learning a sequence of discrete IDs for each item, known as Semantic IDs (Rajput et al., 2023; Singh et al., 2024).
 We explored both approaches and ultimately chose to use dense embedding vectors.
-In our experiments, we observed that Semantic IDs frequently suffered from representational collapse—many items were assigned the same Semantic ID, as also noted in (MultiBehaviorSASRec24; Hourglass24)—and consistently produced lower evaluation metrics (see Table 1(UnifyingGenRetrieval24).
+In our experiments, we observed that Semantic IDs frequently suffered from representational collapse—many items were assigned the same Semantic ID, as also noted in (Liu et al., 2024; Kuai et al., 2024)—and consistently produced lower evaluation metrics (see TIGER in Table 1(Yang et al., 2024).
 While optimizing performance with Semantic IDs remains an interesting direction for future work, our core contributions are representation-agnostic.
@@
 Pin Embeddings.
-For Pins, we combine semantic embedding features with learned Id embeddings (10.1145/3640457.3688053).
-The semantic representation uses the pre-trained OmniSage embedding (badrinath2025omnisage), which captures visual and textual content along with the engagement via the Pin-board graph.
-Additionally, we implement a large-scale ID embedding system where each Pin identifier is hashed into $k$ different hash functions, each mapping to embedding tables of size $d/k$.
-These $k$ embeddings are concatenated to form a $d$-dimensional ID embedding, providing a distributed representation that scales efficiently.
-The ID embeddings constitute sparse parameters that allow us to scale model capacity with minimal impact on serving latency and computational cost, as only a small subset of embeddings are accessed for any given batch of items.
-The concatenated OmniSage and ID embeddings are passed through 2-3 fully connected layers with ReLU activation and layer normalization, followed by $L_{2}$ normalization to produce the final Pin embedding $\mathbf{i}_{\text{pin}}$.
+For Pins, we use the pre-trained OmniSage embedding (Badrinath et al., 2025), which captures visual and textual content along with the engagement via the Pin-board graph.
+The OmniSage embedding is then passed through 2-3 fully connected layers with ReLU activation and layer normalization, followed by $L_{2}$ normalization to produce the final Pin embedding $\mathbf{i}_{\text{pin}}$.
 Query Embeddings.
-For search queries, we use only semantic representations as ID embeddings have not provided value in offline metrics for this entity type.
-We leverage the pre-trained OmniSearchSage query representation (agarwal2024omnisearchsage) which captures semantic query understanding.
+For search queries, we leverage the pre-trained OmniSearchSage query representation (Agarwal et al., 2024) which captures semantic query understanding.
 This representation is passed through a similar MLP architecture (2-3 fully connected layers with ReLU activation and layer normalization) followed by $L_{2}$ normalization to generate the final query embedding $\mathbf{i}_{\text{query}}$.
@@
 For relative time differences, we use log-scale frequencies augmented with learnable phase parameters.
+Our item representations rely on embeddings and signals that are available for new items unseen during training.
+This enables PinRec to generalize to unseen items—such as Ads Pins and Search queries—as we show in Sec.
+6.3, highlighting the strong generalization of generative retrieval models.
 Sequence Construction.
@@
 4.3. Training Objective
-During training, we employ teacher forcing with a next-item prediction objective.
+As described in Sec.
+1, PinRec is trained in two stages: pretraining and fine-tuning.
+Below, we formulate the training objective for each stage.
+Both stages use the same softmax-based loss; they differ only in the target outputs used for supervision and negative examples.
+During both training stage (pretraining and fine-tuning), we employ teacher forcing with a next-item prediction objective.
 The hidden states $\mathbf{h}_{u,t}$ are processed through the conditioned output head $O$ to generate item representations that predict future items in the sequence.
 For outcome-conditioned generation, we condition the output head on the actual values of the target item during training.
-For multi-token prediction, we sample $k$ positions from the next $N$ positions in the sequence to serve as prediction targets.
-Following Pinnerformer22, we compute similarity between the predicted item representation $\hat{\mathbf{i}}_{u,t}$ and candidate item $\mathbf{i}_{c}$ as $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{i}_{c})$, where $Q(\mathbf{i}_{c})$ is a bias correction term estimated using count-min sketches (cormode2005improved) to account for the non-uniform sampling of in-batch negatives.
+Following Pancha et al.
+(2022), we compute similarity between the predicted item representation $\hat{\mathbf{i}}_{u,t}$ and candidate item $\mathbf{i}_{c}$ as $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{i}_{c})$, where $Q(\mathbf{i}_{c})$ is a bias correction term estimated using count-min sketches (Cormode and Muthukrishnan, 2005) to account for the non-uniform sampling of in-batch negatives.
 The sampled softmax loss maximizes similarity to target items relative to negative samples $\mathbf{N}$:
 (2) $L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}})=-\log\frac{\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}}))}{\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{\text{target}}))+\sum_{\mathbf{i}_{n}\in\mathbf{N}}\exp(s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{n}))}$
-We average this loss over all positions in each user sequence. For standard next-token prediction, the target is simply the immediate next item:
-(4) $L_{\theta}^{\text{next}}(H(u,t_{\max}))=\frac{1}{|H(u,t_{\max})|}\sum_{t=1}^{t_{\max}-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
-However, this strict sequential assumption may not align well with user interaction patterns.
-User interaction sequences differ fundamentally from language: while they exhibit global structure reflecting long-term interests, they lack the local structural constraints found in natural language.
-On social media platforms, users engage with feed suggestions in arbitrary order within sessions, making strict positional dependencies less meaningful than in natural language modeling.
-To better accommodate these interaction patterns, we introduce an intra-feed relaxation that considers all items engaged within the same feed session as valid targets.
-Given a feed session $F_{s}=\{i_{t_{1}},i_{t_{2}},\ldots,i_{t_{n}}\}$ where $t_{1}<t_{2}<\ldots<t_{n}$, we define the relaxed target set for position $t_{j}$ as:
-$T_{t_{j}}^{\text{feed}}=\{i_{t_{k}}\mid t_{k}\in F_{s},t_{k}>t_{j}\}$
-This enables a relaxed objective that optimizes for predicting any future item within the same session:
-(5) $L_{\theta}^{\text{feed}}(H(u,t_{\mathrm{max}}))=\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t_{j}}\min_{i\in T_{t_{j}}^{\text{feed}}}L_{s}(\hat{\mathbf{i}}_{u,t_{j}},i)$
-where we minimize the loss to the best-matching item among all future items within the same feed session, rather than strictly requiring prediction of the immediate next item.
-Intuitively, this optimization can be posed as: can we predict any of the future items that the user will engage with within the same feed session?
-We find that this approach is more naturally aligned with user interaction patterns on Pinterest compared to strict next-token prediction.
-Finally, we minimize the expected loss across all user sequences in dataset $\mathcal{D}$:
-(6) $\arg\min_{\theta}\mathbb{E}_{H(u,t_{\max})\sim\mathcal{D}}[L_{\theta}^{\text{feed}}(H(u,t_{\max}))].$
+Pretraining.
+We pretrain the model with next-token prediction on the vast amount of life long user history sequence $H(u,t_{\text{max}})$ (Sec.
+3), which contains actions from across all surfaces in Pinterest and excludes impression-only items (Fig.
+2).
+We compute the loss in Eq.
+2 at each position in the user history sequence and average it across positions.
+(3) $L_{\theta}^{\text{pre}}(H(u,t_{\max}))=\frac{1}{|H(u,t_{\max})|}\sum_{t=1}^{t_{\max}-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
+Fine-tuning.
+For surface-specific fine-tuning, we collect the feed session sequence $F(u,s,t)$ for user $u$ on surface $s$ at time $t$, together with the user’s history up to time $t$ to avoid leakage, $H(u,t)$ (Sec.
+3).
+To perform next-action prediction over both contexts, we construct a combined activity sequence $G$ by concatenating $H(u,t)$ with the action-only entries from $F(u,s,t)$:
+[…]
+We then compute the sequence loss on $G(u,s,t)$ in the same way as in pretraining, averaging the per-position losses.
+(4) $L_{\theta}^{\text{ft}}(G(u,s,t))=\frac{1}{|G(u,s,t)|}\sum_{t=1}^{|G|-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
+Additionally we add impression-only items in $F(u,s,t)$ to the in-batch negative set $\mathbf{N}$ to serve as hard negatives (Fig.
+2).
+We show in Sec.
+5 that these impression-based negatives improve performance.
 4.4. Inference
@@
 The model operates in an autoregressive fashion to generate multiple candidate embeddings, which are then used to retrieve actual items through approximate nearest neighbor search.
-To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer. The specific generation process varies based on the model configuration:
+To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer. The specific generation process varies based on the model configuration (Fig 3):
 Unconditional Generation.
-For the baseline model, generation is straightforward: at each step, the model produces a single item embedding, which is appended to the sequence along with the target surface embedding and a predefined temporal offset (e.g., 10 seconds).
+Without outcome conditioning, generation is straightforward: at each step, the model produces a single item embedding, appended to the sequence along with the target surface embedding and a predefined temporal offset $o=10s$.
 This updated sequence is then used to generate the next embedding.
 Outcome-Conditioned Generation.
-For models with outcome conditioning, for each outcome we generate a candidate embedding, producing multiple candidate embeddings per generation step.
-For autoregressive continuation, we sample a value for each condition according to the requested distribution over each condition $D_{1},D_{2},\ldots,D_{n}$ and use the corresponding embedding as input for the next step (Fig.
-2).
-Temporal Multi-Token Generation.
-When using temporal multi-token prediction, we sample time deltas up to a specified maximum and generate multiple embeddings at different temporal offsets, combined with any specified outcome conditions.
-For the next autoregressive input, we use the embedding corresponding to the immediate next token.
-Across $n$ rollouts, this process generates varying numbers of embeddings depending on the model configuration and input settings.
-The generated embeddings are then used for candidate retrieval through Approximate Nearest Neighbor (ANN) search.
+We generate one candidate embedding per outcome at each step, yielding multiple candidates per generation step.
+The overall FLOPs increase only marginally because a single forward pass can produce embeddings for multiple outcomes, similar to a multitask model.
+For autoregressive continuation, we sample one outcome from the set of considered outcomes and feed its corresponding embedding into the next step (Fig.
+3).
 Budget Allocation.
@@
 The generated candidates can be further refined by downstream ranking models, allowing the system to benefit from both the generative model’s ability to capture complex user patterns and the precision of specialized ranking algorithms.
-The outcome-conditioning and multi-token capabilities make the system particularly suitable for multi-objective optimization scenarios common in production recommendation systems.
+The outcome-conditioning makes the system well suited to the multi-objective optimization needed to customize recommendations for each surface.
 4.5. Serving
@@
 Green boxes represent services, blue rectangles represent transmitted data, purple cylinders represent indexed data, and beige boxes are steps within the PinRec NVIDIA Triton ensemble.
-PinRec is implemented with an in-house, optimized implementation of transformer and served with NVIDIA Triton Inference Server’s ensemble setting and Python backend to perform autoregressive generative retrieval.
-This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with learned ID embeddings, autoregressive inference for item generation, and Faiss retrieval (Figure 3).
+PinRec uses an in-house optimized transformer implementation and served with NVIDIA Triton Inference Server’s ensemble setting and Python backend to perform autoregressive generative retrieval.
+This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with Pin features, autoregressive inference for item generation, and Faiss retrieval (Figure 4).
 4.5.1. Signal Fetching
@@
 The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, storing the sequences and features for each user ID.
-The real-time component is backed by a RocksDB key-value store (kvstore) that efficiently retrieves user engagements occurring after the batch signal’s cutoff time.
+The real-time component is backed by a RocksDB key-value store (Prasad, 2024) that efficiently retrieves user engagements occurring after the batch signal’s cutoff time.
 The batch and real-time signals are deduplicated and returned to PinRec.
@@
 To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision and later dequantized to FP16 for inference.
-4.5.2. Learned ID Embeddings
-As discussed in Section 4.2, our pin embedder includes an ID embedding table that maps each pin ID’s hash to an embedding, followed by MLP layers.
-Since this ID embedder requires significant memory, especially when replicated, it runs as a separate CPU memory-optimized service to avoid competing with the main model for GPU resources.
-It is served with torchscript to achieve single-digit millisecond latencies.
 4.5.2. Autoregressive Inference
@@
 4.5.3. Faiss Retrieval
-(a) Recent user history specified to produce results in Subfigure 4(b).
-(b) Results for the user sequence specified in Subfigure 4(a).
-Figure 4.
-Item recommendations from PinRec-OC for a user history with cat-related search queries and pin interactions (history shown in reverse chronological order).
-Additional examples are provided in Section A.2.
 Our nearest-neighbor retrieval involves a Faiss IVF-HNSW index running on CPU hosts, with inner product as the similarity metric.
@@
 We run kkNN retrieval to fetch the allocated number of Pins for each embedding, with an overfetch factor for later de-duplication.
-Sample results from this complete serving pipeline are visualized for a manually specified user history in Figure 4.
+4.5.4. Serving Efficiency
+On the memory front, PinRec uses about 1.6GiB of GPU memory including KVCache and CUDA graphs, without any significant memory usage increase for outcome conditioning.
+Excluding the KVCache and CUDA graphs, we use roughly the same amount of memory ($\approx$250 MiB) relative to existing two-tower retrieval.
+In terms of latency, at 80 QPS, PinRec-OC’s latency was p50 40ms/p90 65ms, and PinRec-UC’s was p50 23ms and p90 46ms.
+Though the latency is roughly 3-4x higher (p50) than a traditional two-tower solution, we note that parallelizing it with other calls results in ¡1% added end-to-end increase in latency.
 5. Offline Evaluation
-In this section, we dissect the performance of the proposed generative recommendation system, PinRec, on generative retrieval across a range of offline evaluations.
-For the results presented, PinRec uses a 12-layer transformer model (12 attention heads, 768 hidden dimensions) with an 80M-row ID embedding table (256 total dimensions, 8 dimensions per hash).
+In this section, we dissect the performance of PinRec on generative retrieval across a range of offline evaluations.
+For the results presented, PinRec uses a 12-layer transformer model (12 attention heads, 768 hidden dimensions).
 Our experimental setting includes interactions on three major surfaces for the Pinterest platform, namely Home Feed, Related Pins, and Search.
-We demonstrate that outcome-conditioned generation allows for improved controllability and that multi-token generation leads to significantly improved diversity, performance, and efficiency.
+We demonstrate that outcome-conditioned generation captures user intent more accurately, and that surface-specific fine-tuning further improves metrics on each surface.
+Importantly, these two techniques are complementary: neither makes the other redundant.
+Each provides additional gains, and their benefits accumulate when used together.
 5.1. Offline Evaluation Metrics
-We construct the evaluation set using interactions from users disjoint from the training set and evaluate whether the proposed technique can produce representations that retrieve future item embeddings from a set of $1M$ random items (the ”index”).
-For such retrieval tasks, recall@$k$ is a standard performance metric, measuring the proportion of target items correctly retrieved within the top-$k$ ranked items.
+We construct the evaluation set using user activities disjoint from the training set and evaluate whether models can produce representations that retrieve future item embeddings from a set of $1M$ random items (the ”index”).
+We use recall@$k$ (Pancha et al., 2022; Ying et al., 2018; Yang et al., 2024), a standard performance metric measuring the proportion of target items correctly retrieved within the top-$k$ ranked items.
 We evaluate model performance using a sequence generalization of recall@$k$ that accounts for the multi-embedding nature of our autoregressive generation.
-Unlike standard recall, which compares individual predictions to individual targets, our unordered recall metric compares sets of predictions to sets of targets for each user.
+Unlike standard recall, which compares individual predictions to individual targets, our unordered recall metric compares sets of autoregressive generations to sets of future targets for each user.
 Given output item representations $\hat{\mathbf{I}}_{u}=\{\hat{\mathbf{i}}_{u,1},\hat{\mathbf{i}}_{u,2},\ldots,\hat{\mathbf{i}}_{u,m}\}$ and target item embeddings $\mathbf{I}_{u}=\{\mathbf{i}_{u,1},\mathbf{i}_{u,2},\ldots,\mathbf{i}_{u,n}\}$ for user $u$, along with random negative embeddings $\mathbf{N}$, we calculate the proportion of target items for which at least one predicted embedding retrieves it within the top-$k$ results:
@@
 This metric captures whether any of the generated embeddings successfully retrieves each target item, reflecting the practical scenario where multiple predictions can contribute to successful recommendation.
-We perform autoregressive generation without teacher forcing, ensuring the transformer is not conditioned on any target items during prediction, making the evaluation more realistic for deployment scenarios.
-Beyond retrieval performance, we examine recommendation diversity by computing the proportion of unique items retrieved across all users in the evaluation set after selecting top-$k$ predictions.
-This quantifies whether learned representations exhibit bias toward popular items versus providing diverse recommendations across the user base.
+We perform autoregressive generation without teacher forcing, ensuring the transformer is not conditioned on target items during prediction, making evaluation realistic for deployment scenarios.
+Table 1. Comparison of baselines and PinRec variants (unordered recall @ 10) across major surfaces at Pinterest.
+Model Home Feed Related Pins Search
+SASRec (Kang and McAuley, 2018b) 0.382 0.426 0.142
+PinnerFormer (Pancha et al., 2022) 0.461 0.412 0.257
+TIGER (Rajput et al., 2023) 0.208 0.230 0.090
+HSTU (Zhai et al., 2024) 0.596 0.539 0.179
+PinRec-UC 0.608 0.521 0.350
+PinRec-OC 0.625 0.537 0.352
 5.2. Offline Results
-In this section, we compare PinRec against existing sequence-based user modeling approaches, including Pinnerformer (Pinnerformer22), SASRec (kang2018self), and the leading generative recommendation method TIGER (Tiger23).
-We also evaluate our transformer architecture against the proposed HTSU (HSTU24) module.
-Furthermore, we quantify the impacts of our design choices on both recommendation performance and diversity.
-5.2.1. Comparison with Prior Method
-We evaluate three different variants of PinRec.
-  • PinRec-UC: We do not condition on anything (i.e., $\Phi=\emptyset$) and do not employ multi-token prediction.
-  • PinRec-OC: We condition on the target action (“outcome conditioned”) and do not employ multi-token prediction.
-  • PinRec-{MT, OC}: We condition on the target action and employ multi-token prediction.
-To ensure that the comparison between PinRec-{MT, OC} and other methods is fair, we adjust the number of autoregressive generation steps for multi-token generation to ensure that the total number of generated embeddings is held constant.
-As in training (where we cannot modify the evaluation dataset post-hoc), we condition on the target item’s action for outcome-conditioned variants.
-For simplicity, we reproduce results for the outcome-conditioned variant only when comparing against HTSU (additional details in Section A.1.1).
-The impact of negative sampling and model scaling effects are analyzed in Section A.1.2 and Section A.1.3, respectively.
+In this section, we compare PinRec against existing models and quantify the impacts of outcome-conditioning, pretraining and fine-tuning.
+5.2.1. Model Evaluation
+We compare PinRec against existing sequence-based user modeling approaches, including Pinnerformer (Pancha et al., 2022), SASRec (Kang and McAuley, 2018b), and the leading generative recommendation method TIGER (Rajput et al., 2023) and HSTU (Zhai et al., 2024).
+We evaluate different variants of PinRec.
+  • PinRec-UC: We do not condition on anything (i.e., $\Phi=\emptyset$)
+  • PinRec-OC: We condition on the target action (“outcome conditioned”)
+To make it easier and fair to evaluate many model architectures, we use the same training procedure for all methods; all models are trained from the activity sequence from all surfaces like PinRec pretraining (Eq.
+3).
+We evaluate on user activity sequences from held-out users (i.e., users not seen during training).
+There is no fine-tuning process involved in this evaluation.
+Additional details for HSTU evaluation, the impact of negative sampling and model scaling effects are in Sec.
+A.1.1, A.1.2 and A.1.3, respectively.
 Our comparison in Table 1 shows that PinRec outperforms other baselines across the board.
 Importantly, outcome-conditioning demonstrates notable improvements over PinRec-UC (unconditioned) in unordered recall @ 10 (e.g., +2-3% on Home Feed and Related Pins).
-Above all, PinRec-{MT, OC} shows improvements over every method, with over +10% lift on Homefeed, over +20% on Related Pins, and +30% on search over PinRec-UC.
-Table 1. Comparison of baselines and PinRec variants (unordered recall @ 10) across major surfaces at Pinterest.
+5.2.2. Fine-tuning Evaluation
+We evaluate the impact of fine-tuning using surface-view logs.
+Since surface traffic and content evolve over time, we adopt a time split: we train on logs before a cutoff time and evaluate on logs collected after the cutoff, following standard industrial recommender evaluation protocols (Chen et al., 2025; Xia et al., 2023).
+This time-split evaluation set differs from the evaluation data used in Table 1.
+Table 2 shows the impact of pretraining and fine-tuning (FT).
+PinRec-OC (the top row) is the baseline where we test the pretrained model as-is, without no fine-tuning.
+We show the relative lift of other methods, $\frac{\text{method}-PinRec\text{-}OC}{PinRec\text{-}OC}$.
+We evaluate different variants of fine-tuning:
+  • {FT}: Fine-tuning with using surface-specific actions as additional positives (Eq. 4).
+  • {FT, IN} (Impression as Negatives): Impressions only items are used as in-batch negatives during surface fine-tuning.
+  • {FT, IP} (Impression as Positives): Impressions only items in the surface are treated similarly as engaged items.
+We consistently see the strong gains from fine-tuning, supporting our claim that each surface benefits from tuning the model to learn its distinct user behavior patterns and user/item demographics.
+Performance improves further when we add impression-only items as in-batch negatives, since they act as hard “seen-but-not-chosen” counterexamples.
+In contrast, treating impression-only items as positives (IP) rewards exposure rather than intent, introducing noisy positives.
+5.2.3. Outcome-Conditioning’s Impact on Fine-tuning
+We further show that fine-tuning without outcome conditioning remains inferior to fine-tuning with outcome conditioning, supporting our claim that outcome conditioning is essential for fully optimizing GR performance on each surface.
+The last row of Table 2 reports the relative change of the fine-tuned unconditioned model PinRec-UC compared to the pretrained non-fine-tuned outcome-conditioned model PinRec-OC: $\frac{PinRec\text{-}UC\{FT,IN\}-PinRec\text{-}OC}{PinRec\text{-}OC}$.
+For all surfaces, PinRec-UC significantly underperforms PinRec-OC even after fine-tuning, indicating that fine-tuning alone cannot compensate for the absence of outcome conditioning.
+Together, these results show that outcome conditioning is necessary, and that combining it with fine-tuning yields the best performance.
+Table 2. Comparison of fine-tuning PinRec-OC variants (recall @ 10) across major surfaces at Pinterest.
 Model Home Feed Related Pins Search
-SASRec (kang2018self) 0.382 0.426 0.142
-PinnerFormer (Pinnerformer22) 0.461 0.412 0.257
-TIGER (Tiger23) 0.208 0.230 0.090
-HTSU-OC (HSTU24) 0.596 0.539 0.179
-PinRec-UC 0.608 0.521 0.350
-PinRec-OC 0.625 0.537 0.352
-PinRec-{MT, OC} 0.676 0.631 0.450
-5.2.2. Outcome-Conditioned Generation
-We assess whether conditioning on a desired action (e.g., “repin”) leads to outputs that elicit that action.
-If our recommendations successfully control outcomes, we expect higher offline recall for the conditioned action and lower recall for others.
-As a baseline, we compare against the unconditioned PinRec variant.
-For simplicity, we allocate the entire budget to one action and perform offline evaluation conditioning across each action.
-We show the result of this analysis in Figure 5.
-Whenever we condition on an action outcome, we show notable gains over an unconditioned baseline for that action.
-Specifically, matching outcome conditioning yields 1.9% recall lift for true repins, 1.4% on true grid clicks, 4.8% on long grid clicks, and 6.2% on outbound clicks.
-On the other hand, we observe decreases in recall compared to an unconditioned baseline if the true action does not correspond to the desired action, indicating that our conditioning achieves the desired result of producing outcome-specific representations.
-Especially for outbound clicks, we observe that conditioning on any other actions leads to massive drops in unordered recall.
-Figure 5. Percentage lift in unordered recall for PinRec-OC over PinRec-UC when conditioning on the desired action, stratified by the actual action taken by the user.
-5.2.3. Utility of Multi-Token Generation
-We evaluate the benefits of multi-token generation as a function of the number of generations produced per timestep.
-Specifically, we analyze the tradeoff between generating more representations in fewer steps versus fewer representations over more steps, while keeping the total number of output embeddings constant.
-Since the strongest offline performance aside from the multi-token variant is by PinRec-OC, we consider it as a baseline.
-For simplicity, we average the unordered recall @ 10 across the three surfaces.
-We show the tradeoff between performance, diversity, and efficiency with the number of embeddings we generate in each generation step using multi-token generation in Figure 6.
-While next-token generation (i.e., 1 embedding per step) is the most inefficient and yields the poorest performance (where we perform similarly to PinRec-OC), we observe significant benefits across the board in increasing the number of embeddings per step.
-At 16 generations per step, we demonstrate a $\sim$10x reduction in latency over PinRec-OC, with improved unordered recall (+16.0%) and diversity (+21.3%).
-Figure 6. Tradeoff between unordered recall @ 10 (top), proportion unique (bottom), and generation latency (right) with the number of multi-token generations per step.
+PinRec-OC (no FT) 0.0% 0.0% 0.0%
+PinRec-OC {FT} +4.1% +2.1% +2.5%
+PinRec-OC {FT, IN} +4.5% +2.0% +2.5%
+PinRec-OC {FT, IP} +4.1% -1.0% -4.3%
+PinRec-UC {FT, IN} -8.1% -5.5% -9.4%
+Table 3. Percentage lift in recall@10 for PinRec-OC over PinRec-UC when matching outcomes and desired actions.
+Surface Outcome Recall Lift
+Home Feed Save +1.9%
+Outbound Click +6.2%
+Search Product Outbound Click +16%
+All other actions +8%
+5.2.4. Outcome-Conditioning’s Impact on Metrics
+We assess whether conditioning on a desired action (e.g., “save”) leads to outputs that elicit that action.
+If our recommendations successfully control outcomes, we expect higher offline recall for the conditioned action.
+With PinRec-OC{FT, IN} models, we allocate the entire budget to a single action and run recall only for the action (Table 3).
+We see larger improvements for rarer outcomes.
+For example, product outbound click is a rare but important search action, since Pinterest wants users to click through to products they intend to shop.
+In this case, PinRec-OC delivers a 16% lift.
 6. Online A/B Experiments
-We perform online A/B experiments to evaluate PinRec as a candidate generator, measuring its impact on user engagement and confirming its online controllability.
-The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other candidate generators, to produce the final results shown to users.
+We perform online A/B experiments to evaluate PinRec as a Candidate Generator (CG), measuring its impact on user engagement and its ability to tune the trade-off between business metrics.
+The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other Candidate Generators (CGs), to produce the final results shown to users.
 We tested on the three major surfaces: Home Feed, Search and Related Pins.
-6.1. Homefeed
-6.1.1. Experiment Configuration
-In the Homefeed online experiment, we run five groups exposed to global traffic for PinRec-OC, with varying action budgets $\mathcal{B}$.
-The budgets are fixed for outbound click (0.1), long grid clicks (0.05), and other unspecified actions (0.05).
-The remaining budget is split between the repin and grid click engagements, in steps of 0.2 ranging from 0 to 0.8.
-We similarly configure five groups for PinRec-{OC, MT}, with the same budgets.
-We include control groups without PinRec.
+6.1. Trading Off Objectives: A Pareto Analysis
+To validate that PinRec can efficiently control trade-offs between business objectives, we run five PinRec-OC configurations with varying action budgets $\mathcal{B}$.
+We allocate a budget of $\mathcal{B}$ to outcome A and 11-$\mathcal{B}$ to the other outcome.
+Figure 5 shows the resulting Pareto frontier for the two outcome rates (normalized by impressions) produced by the PinRec candidate generator alone; absolute values are omitted due to non-public information.
+These results show that PinRec enables the business to balance multiple metrics by tuning the trade-off between objectives.
+Additional results are shown in Figure 6.
+(a) PinRec-OC Search Goals
+(b) PinRec-OC HF Goals
+Figure 5. Visualization of Pareto fronts for online engagement rates by outcome in surface-specific PinRec variants.
+6.2. Surface A/B Tests
 6.2.1. Metrics
-We use engagement actions such as repins, grid clicks, and time spent.
-Additionally, we count “fulfilled sessions,” where the user takes at least one action in the session, and “unfulfilled sessions,” where the user does not take any action.
-6.1.3. Results
-In Figure 7, we visualize a Pareto front of the repin rate (relative to impressions) and grid click rate for the PinRec candidate generator by itself, with absolute numbers hidden due to non-public information.
-By reducing the grid click budget and correspondingly increasing the repin budget, we see the expected result as the grid click rate drops and repin rate increases.
-Similarly, we observe a significant correlation between the budgets and the overall A/B experiment metric lifts for Homefeed grid clicks and repins in Figure 8 in Appendix.
-(a) PinRec-OC
-(b) PinRec-{MT, OC}
-Figure 7. Visualization of Pareto fronts for online rate of grid click versus repin for items recommended only by PinRec variants, with error bars denoted in grey.
-For each PinRec variant, we selected the variant with the best fulfilled sessions and measured the percentage lift in metrics across Homefeed and the entire site (Table 2).
-Relative to the control group, the unconditioned variant of PinRec, PinRec-UC, demonstrates improvements in unfulfilled sessions and in Homefeed and site-wide grid clicks.
-In addition to those lifts, the outcome-conditioned variant, PinRec-OC, shows significant lift in site-wide fulfilled sessions.
-The multi-token variant, PinRec-{MT, OC}, significantly increased successful site sessions, time spent on site, and site-wide grid clicks.
+We use engagement actions such as saves, grid clicks, and time spent.
+Additionally, we count “fulfilled sessions / fulfillment,” where the user takes at least one action in the session, and “unfulfilled sessions,” where the user does not take any action.
+Note that all reported metric in bold letters are statistically significant at the 5% level (two-sided test; $p<0.05$).
+6.2.2. Home Feed
+For Home Feed, we evaluated both PinRec-UC and PinRec-OC (Table 4).
+While PinRec-UC increases grid clicks, PinRec-OC delivers larger gains in downstream actions and in fulfilled sessions, demonstrating the benefit of outcome-conditioning.
 Table 4. A/B experiments for PinRec variants as candidate generators in Home Feed.
-Type Metric UC OC {MT, OC}
-Overall Fulfilled Sessions +0.02% +0.21% +0.28%
-Metrics Time Spent -0.02% +0.16% +0.55%
-Unfulfilled Sessions -0.28% -0.89% -0.78%
-Action Site-wide Grid Clicks +0.58% +1.76% +1.73%
-Metrics Homefeed Grid Clicks +1.87% +4.01% +3.33%
-6.2. Search
-On the Search surface, we evaluate the unconditioned variant of PinRec.
-As shown in Table 3, PinRec-UC improves search metrics significantly relative to a control group, with increased session-wide search fulfillment (sessions resulting in positive interactions like repins or long clicks).
-When deployed for promoted product retrieval, the same embeddings achieve a 1.83% reduction in CPA and 1.87% increase in shopping conversion volume in the US, indicating improved relevance and user engagement with sponsored content.
+Metric Type Metric UC OC
+Overall Metrics Fulfilled Sessions +0.02% +0.21%
+Action Site-wide Grid Clicks +0.58% +1.76%
+Metrics Home Feed Grid Clicks +1.87% +4.01%
+6.2.3. Related Pins
+Table 5 shows the metric lift from the PinRec-OC on the Related Pins surface, with session-wide gains in fulfilled sessions and sitewide engagement.
+We also see cross-surface impact in Search: fulfillment increases by +0.42% and saves by +0.39%.
+PinRec did not change Search retrieval or ranking; these gains are driven indirectly by better Related Pins recommendations leading to more search actions, highlighting PinRec ’s cross-surface learning.
+Table 5. A/B experiments for PinRec variants as candidate generators in Related Pins.
+Metric Type Metric PinRec-OC
+Overall Fulfilled Sessions +0.21%
+Metrics Time Spent +0.38%
+Action Site-wide Grid Clicks +0.76%
+Metrics Sitewide Saves +0.26%
+6.2.4. Search
+Table 6 shows the results on Search A/B test.
+PinRec-OC improves search metrics significantly relative to a control group, with increased session-wide search fulfillment (sessions resulting in positive interactions like saves or long clicks).
 Table 6. A/B experiments for PinRec as candidate generator in Search retrieval.
-Metric Type Metric PinRec-UC
-Overall Metrics Search Fulfillment Rate +2.27%
-Action Metrics Search Repins +4.21%
-Search Grid Clicks +3.00%
-6.3. Related Pins
-On the Related Pins surface, we evaluate the outcome-conditioned variant of PinRec.
-As shown in Table 4, PinRec-OC improves Related Pins metrics significantly relative to a control group.
-We observe session-wide increase in fulfilled sessions and increase in sitewide engagement.
-Table 4. A/B experiments for PinRec variants as candidate generators in Related Pins.
+Metric Type Metric PinRec-OC
+Overall Metrics Search Fulfillment Rate +2.24%
+Action Metrics Search Saves +3.88%
+Search Shares +5.30%
+6.3. Zero-shot Retrieval
+PinRec represents items and queries using shared, broadly available feature backbones (e.g., OmniSage (Badrinath et al., 2025) and OmniSearchSage (Agarwal et al., 2024)).
+Because these features are defined for a wide range of Pinterest entities, PinRec can be applied to items or queries even when they are absent from the PinRec training corpus.
+We therefore evaluate zero-shot retrieval, asking whether PinRec can retrieve entities that never appear during PinRec training.
+Zero-shot capability is important in practice because Pinterest inventory and user intent evolve continuously: new Pins and ads are created every day, and new or tail queries emerge with limited interaction history.
+A retriever that generalizes zero-shot can serve such cold-start entities immediately, reducing reliance on frequent retraining and improving coverage for long-tail content.
+After running PinRec inference for the PinRec candidate generator (CG), we take the model’s output embedding and perform nearest-neighbor search over other entity corpora:
+  • Related Searches: retrieve search queries whose embeddings are closest to the user’s PinRec embedding.
+  • Ads: retrieve ad Pins using the same nearest-neighbor retrieval approach.
+For both the query and ads corpora, we precompute and index embeddings by running PinRec inference with their respective features described in Sec. 4.2.
+Table 7 reports the results.
+Despite not being trained to predict engagement for these entities, PinRec retrieves highly relevant candidates (e.g., lower cost per ad action).
+Moreover, this retrieval is substantially cheaper in production because PinRec user-side inference is shared with organic CG.
+These results highlight the strong generalization of generative retrieval and support the case for a unified GR model.
+Table 7. A/B experiments for PinRec as a zero-shot candidate generator.
 Type Metric PinRec-OC
-Overall Fulfilled Sessions +0.26%
-Metrics Time Spent +0.30%
-Unfulfilled Sessions -1.07%
-Action Site-wide Grid Clicks +0.50%
-Metrics Sitewide Repins +0.63%
+Related Searches Fulfilled Sessions +1.66%
+(Higher the better) Module Click Rate +4.04%
+Ads Cost Per Acquisition -1.83%
+(Lower the better) Cost Per Click -1.41%
 7. Conclusion
-We introduced PinRec, a generative retrieval system developed for Pinterest’s Homefeed, Search and Related Pins feed.
-To the best of our knowledge, this paper represents the first comprehensive study on deploying generative retrieval in a web-scale application.
-The incorporation of two key innovations — outcome-conditioned generation and temporal multi-token inference — enabled PinRec to enhance several business metrics without substantial cost increases.
+We introduced PinRec, a unified generative retrieval system developed for Pinterest’s Homefeed, Search and Related Pins feed.
+To the best of our knowledge, this paper represents the first comprehensive study on deploying unified generative retrieval in a web-scale application.
+The incorporation of two key innovations — outcome-conditioned generation and surface-specific fine-tuning — enabled PinRec to enhance several business metrics across multiple applications.
 Several promising future directions can be pursued.
 First, applying generative approaches to ranking, as explored by Zhai et al.
-(HSTU24), could be beneficial.
+(Zhai et al., 2024), could be beneficial.
 Second, optimizing input sequences to the model, akin to prompt engineering in large language models, is another avenue worth exploring.
-Third, developing language model alignment techniques similar to Ouyang et al.
-(PPO22) could enhance the use of engagement signals from users.
-Overall, this paper highlights a promising avenue for utilizing generative retrieval in large-scale recommender systems.
+Third, developing language model alignment techniques similar to  He et al.
+(2025) could enhance the use of engagement signals from users.
+Overall, this paper highlights a promising avenue for developing unified generative retrieval for industrial recommender systems.
 Appendix A Appendix
@@
 Though PinRec relies primarily on the well-known transformer architecture, we note that our technique is completely architecture-agnostic.
-Specifically, outcome conditioning and multi-token prediction can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements.
-To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (HSTU24), as opposed to the multi-head attention-based causal transformer.
-Note that we do not leverage the ranking task or any sequence sampling explored in (HSTU24) as it is not relevant for our specific use case.
+Specifically, outcome conditioning can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements.
+To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (Zhai et al., 2024), as opposed to the multi-head attention-based causal transformer.
+Note that we do not leverage the ranking task or any sequence sampling explored in (Zhai et al., 2024) as it is not relevant for our specific use case.
 For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
@@
 A.1.2. Ablating Negative Sampling
-Following from Pinnerformer22, we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function.
-Though we ultimately end up with similar conclusions as Pinnerformer22, we show these findings for maximum clarity regarding our design choices.
+Following from Pancha et al.
+(2022), we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function.
+Though we ultimately end up with similar conclusions as Pancha et al.
+(2022), we show these findings for maximum clarity regarding our design choices.
 Random Negatives
@@
 One of the recent explorations within recommendation systems has been scaling to improve performance, in an attempt to replicate the success of these efforts for language modeling.
-To explore this hypothesis, we scale the number of parameters within PinRec across two fronts: the “vocabulary” and the model itself.
-While model scaling is clear and involves scaling up the embedding size, the number of transformer layers, and so on, scaling the vocabulary for an embedding-based generative retrieval technique involves improving the input representations.
-Specifically, as mentioned in the main text, we implement “ID” embeddings for pins, that effectively serve as additional representation for the “pin vocabulary”.
-Model Parameters
+To explore this hypothesis, we scale the number of parameters within PinRec.
 As presented in the main text, PinRec (all variants) have roughly 100 million transformer parameters, corresponding to the base form or size of GPT-2.
@@
 Though we observe some gains in Table 11 from scaling the number of transformer parameters to around 1 billion (XL-sized), these become intractable to serve in real time, requiring offline batch inference or significantly higher cost and more optimization.
-For the remaining sizes, we observe insignificant lifts (note that less than 1% in lift magnitude can be often explained through noise).
+For the remaining sizes, there was no improvement observed actually.
 Hence, we opt to not use larger sized transformers.
-Vocabulary Parameters.
-As presented in the main text, the ID embeddings consist of roughly 10 billion parameters.
-These are a core component of the model since they provide additional representational power for the most important entity type at Pinterest: Pins.
-To ablate the importance of these ID embeddings, we examine the performance in offline recall across all surfaces with and without ID embeddings.
-For this analysis, we focus on Homefeed (the most critical surface) and present results aggregated across all surfaces as well in Table 9.
-Table 9. Percentage lift through adding ID embeddings (10 billion parameters) over no ID embedding-based vocabulary.
-Surface % Lift in Recall @ 10
-Homefeed +14.0%
-All +14.4%
-Evidently, the improvement in recall is substantial and indicates that the scaling of vocabulary to learn an embedding for each unique ID (roughly, as it is hash-based) rather than only content-based features is beneficial.
 A.1.4. Tail User and Item Analysis
@@
 Table 12. Online experimental results for different user cohorts and tail item metrics, across PinRec-UC and PinRec-OC.
-In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 10(a).
+In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 12(a).
 Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users.
@@
 (a) PinRec-OC
-(b) PinRec-{MT, OC}
-Figure 8. Visualization of Pareto fronts for online lift % in Homefeed grid clicks versus repins for PinRec variants.
-A.1.5. Serving Efficiency
-One of the critical aspects of our proposed generative retrieval technique is an effort to maintain serving efficiency.
-With our custom infrastructure built from off-the-shelf serving technologies such as NVIDIA Triton Inference Server and optimization techniques like torch.compile, we achieve a system that can be served in real-time.
-In this section, we delve deeper into the memory and time requirements, comparing our method with prior methods deployed in production.
-On the memory front, all variants of PinRec (unconditioned, outcome-conditioned with or without multi-token generation) use about 1.6GiB of GPU memory including KVCache and CUDA graphs, without any significant memory usage increase for multi-token prediction or outcome conditioning.
-Excluding the KVCache and CUDA graphs, we use roughly the same amount of memory (2̃50 MiB) relative to other techniques (i.e., two-tower models).
-In terms of latency, at 80 QPS, the online and offline latency for PinRec-OC/OC, MT are p50 40ms/p90 65ms (6 steps).
-On the other hand, across only 3 steps for PinRec-UC, we observe a latency of p50 23ms and p90 46ms (note that this produces significantly less embeddings and less compressed embeddings than either of the above techniques, though it is lower latency).
-Though the latency is roughly 3-4x higher (p50) than a traditional two-tower solution, we note that parallelizing it with other RPCs results in ¡1% added end-to-end increase in latency.
-A.2. Visualization of Multi-Token Impact
-(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
-(b) Results for PinRec-OC candidate generator for user history in subfigure above. Notably only includes Pins related to nails, the user’s most recent interaction.
-(c) Results for PinRec-{MT, OC} candidate generator for user history in subfigure above. Includes a mix of Pins related to both tattoos and nails.
-Figure 9. Comparison of results for a manually-specified user history.
-(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
-(b) Results for PinRec-OC candidate generator for given user history. Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
-(c) Results for PinRec-{MT, OC} candidate generator for given user history. Includes a mix of Pins related to nurseries and food.
-Figure 10. Comparison of results for another manually-specified user history.
-In this section, we present additional qualitative examples through visualizations of retrieved examples for sample user histories, comparing multi-token and non-multi-token variants.
-Through these visualizations, we are able to demonstrate a reasonable and consistent qualitative difference in the types of pins retrieved, justifying the improved session-wide and sitewide online metrics for PinRec-{MT, OC}.
-Given a user history, we display two examples of retrieved pins in Figure 9 and Figure 10 for PinRec-{MT, OC} and PinRec-{OC}.
-Specifically, note that PinRec-{MT, OC} captures less-recent interactions for users compared to PinRec-OC.
-This is expected as the multi-token model predicts over a longer future time window and is less biased toward recency.
-PinRec-{MT, OC} can therefore produce more diverse results capturing more medium-term interests.
-On the other hand, since the PinRec-{OC} is naturally biased towards the short term, it fails to capture potential longer term trends that may induce engagement.
-Note that these trends reflect the same takeaways from offline and online evaluation.
-Offline, we observe an improvement in diversity as the number of multi-token predictions is increased, which is consistently above the non-multi-token variant of PinRec.
-Online, we see that the “feed” presented by PinRec-{MT, OC} is of higher quality due to its diversity and visual appeal/aesthetic.
-These are critical qualities for a platform like Pinterest.
+Figure 6. Visualization of Pareto fronts for online lift % in Home Feed grid clicks versus repins for PinRec
+A.1.5. Embedding similarity threshold tuning
+For the embedding similarity threshold for embedding compression in Sec.
+4.4), we A/B tested values in the [0.75, 0.9] range for each surface.
+Our findings were: (1) results were qualitatively consistent across the different values, and (2) a threshold of 0.9 slightly outperformed other values for both Search and Related Pins, as reported in the paper.
+We will clarify this tuning procedure in the final version.
+We have not yet tuned this value for different user segments, but agree that this would be an interesting future research.
+A.1.6. Temporal offset value tuning
+Inference: We experimented with different values of $\delta$ for each action at inference time.
+For the results reported in this paper, we sampled repins and grid-clicks at $\delta\in\{0,25,50\}$ (i.e., 0, 25, and 50 discrete timesteps ahead).
+For all other outcomes, we sampled only the immediate next timestep ($\delta=0$).
+We also tried sampling every 5 timesteps from 0 to 50, but this generated too many embeddings for ANN search.
+Training: We randomly selected 4 training targets with $\delta$ sampled from the interval $[0,25)$.
+For evaluation, we compared the model outputs at $\delta\in\{0,10\}$ and measured recall@10 for predicting the ground-truth action under different Timestep Gaps (TG).
+TG = 0 corresponds to immediate-term recall, while TG = 30 reflects long-term recall.
+Compared to $\delta=0$, $\delta=10$ achieved 60% higher long-term recall with only a 2% decrease in short-term recall, indicating a better overall trade-off; therefore, we use $\delta=10$.